### v1

# Stabilization of intersection Betti numbers for moduli spaces of one-dimensional sheaves on surfaces

Abstract.
In this paper, we study the intersection cohomology of the moduli space of semistable one-dimensional sheaves with fixed Euler characteristic, supported in a divisor class $\beta$ on a smooth projective surface $S$. Assuming that this moduli space is irreducible, we prove that its intersection Betti numbers in a certain range are determined by a product formula derived from Göttsche’s formula for Betti numbers of Hilbert schemes of points on $S$. As an application, for Enriques and bielliptic surfaces, we show the stabilization of intersection Betti numbers of this moduli space when $\beta$ is sufficiently positive. In the Enriques case, we also prove a refined stabilization result concerning the perverse Hodge numbers when the moduli space is smooth.
1. Introduction
Throughout, we work over the field $\mathbb{C}$ of complex numbers. Let $S$ be a connected, smooth, projective surface with a fixed ample divisor $H$.
In recent years, moduli spaces of one-dimensional sheaves on $S$ have attracted considerable attention. On the one hand, their connections with enumerative geometry and moduli spaces of Higgs bundles have led to a number of new conjectures and results; see, for instance, [3, 4, 9, 14, 15, 19, 21, 28, 29, 36]. On the other hand, these moduli spaces exhibit not only asymptotic behavior analogous to that of moduli spaces of torsion-free sheaves, but also pathologies [34, §5] in contrast to the torsion-free case, revealing rich geometry that remains to be understood.
Let $\chi$ be an integer and $\beta$ a nonzero effective divisor of arithmetic genus $p_{a}(\beta)$ on $S$. Denote by $|\beta|^{\mathrm{int}}\subset|\beta|$ the open subscheme parametrizing integral curves. Consider the (coarse) moduli space $M_{\beta,\chi}$ parametrizing S-equivalence classes of one-dimensional semistable (with respect to $H$) sheaves $\mathcal{F}$} on $S$ with fixed determinant $\operatorname{det}({\mathcal{F}})\cong{\mathcal{O}}_{S}(\beta)$ and Euler characteristic $\chi({\mathcal{F}})=\chi$. Here the stability is defined in terms of the slope
$\mu({\mathcal{F}})=\frac{\chi({\mathcal{F}})}{c_{1}({\mathcal{F}})\cdot H}.$
There is a morphism
$h:M_{\beta,\chi}\to|\beta|,$
called the Hilbert–Chow morphism, which is defined by taking the Fitting support. If $|\beta|$ contains a connected smooth curve, then the morphism $h$ is an abelian fibration.
The moduli space $M_{\beta,\chi}$ is singular in general, for which the intersection cohomology $\operatorname{IH}^{*}(M_{\beta,\chi})$ (with $\mathbb{Q}$}-coefficients) might be a more suitable topological invariant than the ordinary cohomology. In this paper, we prove the following theorem about the intersection Betti numbers of $M_{\beta,\chi}$.
Theorem 1.1.
Fix $k\in\mathbb{Z}_{\geq 0}$ and $\chi\in\mathbb{Z}$. Suppose that $M_{\beta,\chi}$ is irreducible, and let
$c\vcentcolon=\operatorname{codim}(M_{\beta,\chi}\setminus h^{-1}(|\beta|^{\mathrm{int}}),M_{\beta,\chi})$
denote the codimension of $M_{\beta,\chi}\setminus h^{-1}(|\beta|^{\mathrm{int}})$ in $M_{\beta,\chi}$. If the following conditions hold:
  1. (1)
    $\beta$ is $\max\{2,2k-2\}$-very ample (see Definition 2.2);
  2. (2)
    $k\leq\min\{c-1,\frac{2}{3}\dim|\beta|\}$,
then we have
$\dim_{\mathbb{Q}}\operatorname{IH}^{k}(M_{\beta,\chi})=b_{k}^{\infty},$
where $b_{k}^{\infty}$ is defined by the following identity of formal power series
(1.1) $\sum_{k\geq 0}b_{k}^{\infty}q^{k}=\prod_{m\geq 1}\frac{(1+q^{2m-1})^{b_{1}(S)}(1+q^{2m+1})^{b_{1}(S)}}{(1-q^{2m})^{b_{2}(S)+1}(1-q^{2m+2})}.$
By Göttsche’s formula [8], the number $b_{k}^{\infty}$ is exactly the $k$-th Betti number $b_{k}(S^{[n]})$ of the Hilbert scheme $S^{[n]}$ of $n$ points for any integer $n\geq k$. The connection between Betti numbers of moduli spaces of sheaves on $S$ and $b_{k}^{\infty}$ has been observed in many studies, for example, in [5, 17] for some moduli spaces of torsion-free sheaves, in [36] for moduli spaces of one-dimensional sheaves on $\mathbb{P}^{2}$, and in [29] for those of one-dimensional sheaves on any del Pezzo surface.
Our proof of Theorem 1.1 adapts the approach in §2 and §3 of our previous joint work with Pi and Shen [29] for a more general setting. A key new ingredient in the proof is as follows.
Theorem 1.2.
Let $\ell\geq 2$ be an integer. If $\beta$ is $(2\ell-2)$-very ample, then there exists an open subscheme $U_{\ell}\subset|\beta|^{\mathrm{int}}$ such that
  1. (1)
    $\operatorname{codim}(|\beta|^{\mathrm{int}}\setminus U_{\ell},|\beta|^{\mathrm{int}})\geq\ell+1$;
  2. (2)
    The open subscheme $h^{-1}(U_{\ell})$ of $M_{\beta,\chi}$ is smooth.
For simplicity of exposition, now we set $\chi=1$ and consider $M_{\beta}\vcentcolon=M_{\beta,1}$ (see Remark 5.4 for the case of arbitrary $\chi$). As an application of Theorem 1.1, we prove the stabilization of intersection Betti numbers of $M_{\beta}$ for Enriques and bielliptic surfaces, providing new evidence for [33, Conjecture 1.3].
Theorem 1.3 (Theorem 5.3).
Suppose that $S$ is an Enriques or bielliptic surface. Fix $k\in\mathbb{Z}_{\geq 0}$. Given any ample divisor $\beta_{0}$ on $S$, there exists $d(\beta_{0},k)\in\mathbb{Z}_{>0}$ (depending on $\beta_{0}$ and $k$) such that
$\dim_{\mathbb{Q}}\operatorname{IH}^{k}(M_{d\beta_{0}})=b_{k}^{\infty}\quad\text{for all integer }d\geq d(\beta_{0},k),$
where $b_{k}^{\infty}$ is defined by (1.1).
Although moduli spaces of sheaves on $K3$ and abelian surfaces have been extensively studied, much less is known about the cases of Enriques and bielliptic surfaces. In the bielliptic case, recent progress was made in [27].
When $S$ is a generic Enriques surface and the class of $\beta$ is not divisible by $2$ in the group $\operatorname{Num}(S)$ of numerical equivalence classes of divisors on $S$, Saccà [30] proved that $M_{\beta}$ is a smooth projective Calabi–Yau variety of dimension $\beta^{2}+1$. In this case, although the odd cohomology of $M_{\beta}$ can be nonzero as shown in [30, Theorem 5.10], Theorem 1.3 implies that the $k$-th Betti number $b_{k}(M_{\beta})$ for fixed odd $k$ vanishes for sufficiently positive $\beta$. We also prove, in this setting, a refined version of Theorem 1.3 concerning the stabilization of the invariants $n_{\beta}^{i,j}$ (defined by (2.4) with $\phi$ taken to be $h:M_{\beta}\to|\beta|$) which are conjectured to refine the Gromov–Witten/Pandharipande–Thomas invariants of the local surface $\operatorname{Tot}(K_{S})$ [10, 12, 21] and are the shifted perverse Hodge numbers considered in [28].
Theorem 1.4 (Theorem 5.7).
Suppose that $S$ is a generic Enriques surface. Fix $i,j\in\mathbb{Z}_{\geq 0}$. Given any ample divisor $\beta_{0}$ on $S$ such that $2\nmid\beta_{0}$ in $\operatorname{Num}(S)$, there exists $d(\beta_{0},i,j)\in\mathbb{Z}_{>0}$ (depending on $\beta_{0}$, $i$ and $j$) such that
$n_{d\beta_{0}}^{i,j}=n_{\infty}^{i,j}\quad\text{for all integer $d\geq d(\beta_{0},i,j)$ with }2\nmid d,$
where $n_{\infty}^{i,j}$ is defined by the following identity of formal power series
(1.2) $\sum_{i,j\geq 0}n_{\infty}^{i,j}q^{i}t^{j}=(1-qt)\prod_{m\geq 1}\frac{1}{(1-q^{m+1}t^{m-1})(1-q^{m}t^{m})^{10}(1-q^{m-1}t^{m+1})}.$
Remark 1.5.
Under the assumption of [28, Conjecture A], Oberdieck proved a stronger stabilization result [28, Proposition 1.1] for the shifted perverse Hodge numbers. We note that the product formula (1.2) coincides with the one in [28, Proposition 1.1]. It remains open whether our approach, which does not rely on [28, Conjecture A], can be modified to show $n_{\beta}^{i,j}=n^{i,j}_{\infty}$ if $i<\beta^{2}/4$ and $j<\beta^{2}/4-1$. An explicit expression for $d(\beta_{0},i,j)$ in Theorem 1.4 is given in (5.27).
The rest of this paper is organized as follows. In Section 2, we collect some notation, definitions, and facts that will be useful later. Section 3 is devoted to the local versality of the universal family of curves in high codimension and thus to a proof of Theorem 1.2. In Section 4, we prove Theorem 1.1 and discuss some difficulties arising in the general case. Finally, in Section 5, we study in detail the cases of Enriques and bielliptic surfaces and prove Theorems 1.3 and 1.4.
Conventions
A variety is an integral and separated scheme of finite type. A curve is a projective scheme of pure dimension $1$. By a point of a scheme, we always mean a closed point. All sheaves are assumed to be coherent. For a sheaf $\mathcal{E}$} on a projective scheme $X$, we denote by $h^{i}({\mathcal{E}})$ (resp. $\operatorname{ext}^{i}({\mathcal{E}},{\mathcal{E}})$) the dimension of the $i$-th sheaf cohomology $H^{i}(X,{\mathcal{E}})$ (resp. $i$-th Ext group $\operatorname{Ext}^{i}({\mathcal{E}},{\mathcal{E}})$) as a $\mathbb{C}$-vector space.
Acknowledgments
The authors thank Weite Pi and Junliang Shen for helpful discussions. The second author is grateful to her advisor Jun Li for his constant support, to János Kollár and Claire Voisin for inspiring discussions on the deformation theory of singularities of planar curves, and to Georg Oberdieck for suggesting considering bielliptic surfaces. The first author is partially supported by the Fundamental Research Funds for the Central Universities. The second author is supported by the NSFC grants (No. 12121001 and No. 12425105) and LMNS.
2. Preliminaries
This section gathers the notation and background material that will be essential later.
2.1. Relative Hilbert schemes of points
Let $\pi:{\mathcal{C}}\to|\beta|$ be the universal family of curves in $|\beta|$. For $n\in\mathbb{Z}_{\geq 0}$, denote by
$\pi^{[n]}:{\mathcal{C}}^{[n]}\to|\beta|$
the relative Hilbert scheme of $n$ points on the fibers of $\pi$. For $C\in|\beta|$, the fiber of $\pi^{[n]}$ over $C$ is the Hilbert scheme $C^{[n]}$ parametrizing 0-dimensional, length $n$ closed subschemes of $C$. Note that $\pi^{[0]}:|\beta|\to|\beta|$ is the identity and $\pi=\pi^{[1]}:{\mathcal{C}}\to|\beta|$. For any open subscheme $U\subset|\beta|$, write ${\mathcal{C}}_{U}^{[n]}$ for $(\pi^{[n]})^{-1}(U)$. The following property of $\pi^{[n]}$ is a corollary of [18, Theorem 1.1] (see also [34, Proposition 2.7]).
Proposition 2.1.
The morphism $\pi^{[n]}:{\mathcal{C}}^{[n]}\to|\beta|$ has fibers of the same dimension $n$.
There is a natural morphism $\sigma_{n}:{\mathcal{C}}^{[n+1]}\to S^{[n+1]}$ defined by
(2.3) $Z\subset C]\mapsto[Z\subset S],$
where $C\in|\beta|$ and $Z\subset C$ is a 0-dimensional closed subscheme of length $n+1$. The fiber of $\sigma_{n}$ over $Z\subset S]\in S^{[n+1]}$ is the projective space $\mathbb{P}(H^{0}(S,{\mathcal{I}}_{Z}\otimes{\mathcal{O}}_{S}(\beta))^{\vee})$, where ${\mathcal{I}}_{Z}$ is the ideal sheaf of $Z\subset S$.
2.2. Positivity conditions
We recall the following notion of $k$-very ampleness, which generalizes base-point-freeness (0-very ampleness) and very ampleness ($1$-very ampleness).
Definition 2.2.
An effective divisor $\beta$ on $S$ is called $k$-very ample $(k\in\mathbb{Z}_{\geq 0})$ if for any 0-dimensional closed subscheme $Z\subset S$ of length $k+1$, the restriction map
$r_{Z}:H^{0}(S,{\mathcal{O}}_{S}(\beta))\to H^{0}(Z,{\mathcal{O}}_{S}(\beta)|_{Z})$
is surjective. We adopt the convention that every divisor is (1)(-1)-very ample.
The definition immediately implies two properties, which will be used in the sequel:
  1. (1)
    If $\beta$ is $k$-very ample and $k^{\prime}\leq k$ $(k^{\prime}\in\mathbb{Z}_{\geq 0})$, then $\beta$ is also $k^{\prime}$-very ample.
  2. (2)
    If $\beta$ is ample, then for any $k\in\mathbb{Z}_{\geq 0}$, there exists $d_{k}\in\mathbb{Z}_{>0}$ such that the multiple $d\beta$ is $k$-very ample for every integer $d\geq d_{k}$.
A well-known connection between $k$-very ampleness and the structure of ${\mathcal{C}}^{[k+1]}$ as a scheme over $S^{[k+1]}$ via $\sigma_{k}$ is as follows (e.g., [34, Proposition 2.6]).
Proposition 2.3.
If $\beta$ is $k$-very ample, then ${\mathcal{C}}^{[k+1]}$ is a projective bundle over $S^{[k+1]}$.
2.3. Perverse filtrations
Next we recall some generalities of perverse filtrations. Let $\phi:X\to Y$ be a proper morphism between smooth quasiprojective varieties. We assume that $\dim X=a$, $\dim Y=b$, and all fibers of $\phi$ have dimension aba-b. The perverse filtration
$P_{0}H^{m}(X)\subset P_{1}H^{m}(X)\subset\dots\subset H^{m}(X)$
induced by $\phi$ is an increasing filtration on the rational cohomology $H^{*}(X)$, governed by the topology of the morphism $\phi$. It is defined to be
$P_{k}H^{m}(X)\vcentcolon=\mathrm{Im}\left\{H^{m-b}(Y,{{}^{\mathbf{p}}\tau_{\leq k}}(R\phi_{*}{\mathbb{Q}}_{X}[b]))\to H^{m-b}(Y,R\phi_{*}{\mathbb{Q}}_{X}[b])\right\}\subset H^{m}(X),$
where $^{\mathbf{p}}\tau_{\leq\bullet}$ is the perverse truncation functor [2]. If we apply the decomposition theorem to $\phi:X\to Y$, we obtain
$R\phi_{*}{\mathbb{Q}}_{X}[b]\simeq\bigoplus_{i=0}^{2a-2b}\mathcal{P}_{i}[-i]\in D^{b}_{c}(Y)$
with $\mathcal{P}_{i}$ a semisimple perverse sheaf on $Y$. There is an identity
$P_{k}H^{m}(X)=\mathrm{Im}\Big\{H^{m-b}(Y,\bigoplus_{i=0}^{k}\mathcal{P}_{i}[-i])\to H^{m}(X)\Big\}.$
The dimension of the graded piece $\mathrm{Gr}_{i}^{P}H^{i+j}(X)\vcentcolon=P_{i}H^{i+j}(X)/P_{i-1}H^{i+j}(X)$ is denoted by
(2.4) $n_{\phi}^{i,j}\vcentcolon=\dim_{\mathbb{Q}}\mathrm{Gr}_{i}^{P}H^{i+j}(X)=\dim_{\mathbb{Q}}H^{j-b}(Y,{\mathcal{P}}_{i}).$
Note that for the Hilbert–Chow morphism $h:M_{\beta,\chi}\to|\beta|$, the fibers of $h$ over integral curves in $|\beta|$ are the compactified Jacobians of degree $p_{a}(\beta)-1+\chi$ and have the same dimension $p_{a}(\beta)$. Later, we will use the following result of Maulik–Yun [22, Theorem 2.13] and Migliorini–Shende [24, Corollary 2].
Theorem 2.4 ([22, 24] [^1] The result of [22, 24] can be applied here since the smoothness of $h^{-1}(U)$ implies that all relative compactified Jacobians $\overline{J}_{U}^{e}$ ($e\in\mathbb{Z}$) are smooth by the existence of a section of the universal curve (étale locally on $U$, avoiding the singularities of fibers) and so are all relative Hilbert schemes ${\mathcal{C}}^{[\ell]}_{U}$ ($\ell\in\mathbb{Z}_{\geq 0}$) by [32, Proposition 14].).
Suppose that $|\beta|$ contains a connected smooth curve. Let $U\subset|\beta|^{\mathrm{int}}$ be an open subscheme such that $h^{-1}(U)$ is smooth, and let $P_{\bullet}H^{*}(h^{-1}(U))$ be the perverse filtration induced by the restriction $h_{U}:h^{-1}(U)\to U$ of $h$. For any $\ell,m\in\mathbb{Z}_{\geq 0}$, we have an isomorphism of vector spaces
(2.5) $H^{m}\left({\mathcal{C}}^{[\ell]}_{U}\right)\cong\sum_{i+j\leq\ell;\,i,j\geq 0}\mathrm{Gr}_{i}^{P}H^{m-2j}\left(h^{-1}(U)\right).$
Remark 2.5.
When $\ell=0$, the isomorphism (2.5) reads
$H^{m}(U)\cong\mathrm{Gr}_{0}^{P}H^{m}\left(h^{-1}(U)\right).$
This isomorphism can be obtained as a consequence of Ngô’s support theorem [26, Théorème 7.2.1]. This theorem shows that $h_{U}:h^{-1}(U)\to U$ has full support; see [22, Theorem 2.4, §2.5] or [24, Corollary 9]. More precisely, there is an isomorphism
$Rh_{U*}{\mathbb{Q}}_{h^{-1}(U)}[\dim U]\cong\bigoplus_{i=0}^{2p_{a}(\beta)}\operatorname{IC}(\wedge^{i}R^{1}\pi^{s}_{*}{\mathbb{Q}}_{{\mathcal{C}}^{\mathrm{sm}}})[-i],$
where $\pi^{s}:{\mathcal{C}}^{\mathrm{sm}}\to|\beta|^{\mathrm{sm}}$ is the universal family of smooth curves in $|\beta|$.
3. Singular locus of relative compactified Jacobians
The goal of this section is to prove Theorem 1.2 about the smoothness of a big open subscheme of $h^{-1}(|\beta|^{\mathrm{int}})$. We begin by reviewing some basic notions of plane curve singularities and deformation theory.
3.1. Singularities and deformations of planar curves
3.1.1. Singularity invariants
We first recall some basic invariants of isolated singularities in planar curves.
Definition 3.1.
Let $f\in\mathbb{C}[[x,y]]$ be a convergent power series.
  1. (1)
    The numbers μ(f):=dim_CC[[x,y]]/(∂_x f,∂_y f), τ(f):=dim_CC[[x,y]]/(f,∂_x f,∂_y f) are called the Milnor number and Tjurina number of $f$, respectively.
  2. (2)
    If $f$ is reduced, let O=C[[x,y]]/(f)↪O be the normalization of $\mathcal{O}$}. Then the number δ(f):=dim_CO/O is called the $\delta$-invariant of $f$.
The Milnor number and the $\delta$-invariant are related by the following formula of Milnor.
Proposition 3.2 ([25, §10]).
Let $f\in(x,y)\subset\mathbb{C}[[x,y]]$ be a reduced, convergent power series. Denote by $r(f)$ the number of irreducible factors of $f$. Then
(3.6) $\mu(f)=2\delta(f)-r(f)+1.$
If $C$ is a reduced curve on $S$, then the singular locus $C_{\mathrm{sing}}$ of $C$ is discrete. For any point $p\in C_{\mathrm{sing}}$, there exists a polynomial $f\in(x,y)\subset\mathbb{C}[x,y]$ without multiple factors such that the singularity of $C$ at $p$ is analytically isomorphic to that of $\{f=0\}$ at $(0,0)$, i.e.,
$\hat{\mathcal{O}}_{C,p}\cong\mathbb{C}[[x,y]]/(f).$
Thus, the Milnor and Tjurina numbers, and the $\delta$-invariant of $C$ at $p$ can be defined as those of $f$ in Definition 3.1, denoted by $\mu_{p}(C)$, $\tau_{p}(C)$, and $\delta_{p}(C)$, respectively.
For completeness, we give a proof of the following standard interpretation of the total Tjurina number of $C$ in terms of the Ext sheaf ${\mathcal{E}xt}^{1}(\Omega_{C},{\mathcal{O}}_{C})$, where $\Omega_{C}$ is the cotangent sheaf of $C$.
Lemma 3.3.
If $C$ is a reduced curve on S, then
$\sum_{p\in C_{\mathrm{sing}}}\tau_{p}(C)=h^{0}({\mathcal{E}xt}^{1}(\Omega_{C},{\mathcal{O}}_{C})).$
Proof.
The sheaf $\Omega_{C}$ is locally free over the smooth locus $C\setminus C_{\mathrm{sing}}$ of $C$, so ${\mathcal{E}xt}^{1}(\Omega_{C},{\mathcal{O}}_{C})$ is a torsion sheaf supported on $C_{\mathrm{sing}}$. Let $p\in C_{\mathrm{sing}}$ and assume that the singularity of $C$ at $p$ is analytically isomorphic to that of $\{f=0\}$ at $(0,0)$ for $f\in(x,y)\subset\mathbb{C}[x,y]$ without multiple factors. The lemma can be proved locally around $p$. For $R\vcentcolon=\mathbb{C}[x,y]/(f)$, we have $\Omega_{R}=(Rdx\oplus Rdy)/(\partial_{x}fdx+\partial_{y}fdy)$. By the conormal exact sequence
$0\to(f)/(f^{2})\to R^{\oplus 2}\to\Omega_{R}\to 0,$
we conclude that $\operatorname{Ext}^{1}(\Omega_{R},R)=R/(\partial_{x}f,\partial_{y}f)$, from which the result follows. ∎
3.1.2. Smoothness from local versality
Let $\pi_{B}:{\mathcal{C}}_{B}\to B$ be a family of reduced planar curves, i.e., a proper and flat morphism whose fibers are reduced curves with at worst planar singularities. Assume that $B$ is a smooth variety. There is a natural map, called the local Kodaira–Spencer map:
(3.7) $k_{\pi_{B},b}^{\text{loc}}:T_{b}B\to\prod_{p\in({\mathcal{C}}_{b})_{\mathrm{sing}}}T\mathrm{Def}_{{\mathcal{C}}_{b},p}=H^{0}({\mathcal{C}}_{b},{\mathcal{E}xt}^{1}(\Omega_{{\mathcal{C}}_{b}},{\mathcal{O}}_{{\mathcal{C}}_{b}})),$
where $({\mathcal{C}}_{b})_{\mathrm{sing}}$ is the singular locus of ${\mathcal{C}}_{b}\vcentcolon=\pi_{B}^{-1}(b)$, and $T\mathrm{Def}_{{\mathcal{C}}_{b},p}$ is the tangent space of the deformation functor of the complete local $\mathbb{C}$-algebra $\hat{{\mathcal{O}}}_{{\mathcal{C}}_{b},p}$.
Definition 3.4.
The family $\pi_{B}:{\mathcal{C}}_{B}\to B$ is called locally versal if for any point $b\in B$, the local Kodaira–Spencer map $k_{\pi_{B},b}^{\text{loc}}$ in (3.7) is surjective.
If $\pi_{B}:{\mathcal{C}}_{B}\to B$ is a family of integral planar curves, then for any $e\in\mathbb{Z}$, denote by
$p_{B,e}:\overline{J}_{B}^{e}\to B$
the relative compactified Jacobian of degree $e$ associated with $\pi_{B}$, whose fiber over a point $b\in B$ is the moduli space of rank $1$, torsion-free sheaves of degree $e$ (or equivalently, Euler characteristic $e+1-p_{a}({\mathcal{C}}_{b})$) on ${\mathcal{C}}_{b}$. Since $\pi_{B}$ admits, étale locally on $B$, a section which avoids the singularities of fibers of $\pi_{B}$, the smoothness of $\overline{J}_{B}^{e}$ for any $e$ is equivalent to the smoothness of $\overline{J}_{B}^{0}$. Hence, the following result follows directly from [7, Corollary B.3].
Theorem 3.5 ([7]).
If $\pi_{B}:{\mathcal{C}}_{B}\to B$ is a locally versal family of integral planar curves, then $\overline{J}_{B}^{e}$ is smooth for any $e\in\mathbb{Z}$.
3.2. Local versality and smoothness in high codimension
Now we show that the universal family $\pi:{\mathcal{C}}\to|\beta|$ of curves in $|\beta|$ is locally versal in sufficiently high codimension if $\beta$ is positive enough.
Proposition 3.6.
Let $\ell\geq 2$ be an integer. If $\beta$ is $(2\ell-2)$-very ample, then there exists an open subscheme $V_{\ell}\subset|\beta|$ such that
  1. (1)
    Every point of $V_{\ell}$ corresponds to a reduced curve;
  2. (2)
    $\operatorname{codim}(|\beta|\setminus V_{\ell},|\beta|)\geq\ell+1$;
  3. (3)
    The universal family $\pi|_{\pi^{-1}(V_{\ell})}:\pi^{-1}(V_{\ell})\to V_{\ell}$ of curves in $V_{\ell}$ is locally versal.
Proof.
We use an argument in [16] to control the locus of curves that are not $\ell$-nodal. Here, a curve is called $\ell$-nodal if it is reduced and has no singularities other than $\ell$ nodes. Since $\beta$ is also $\ell$-very ample, by the proof of [16, Proposition 2.1], there is an open subscheme $V_{\ell}\subset|\beta|$ such that
  • $\operatorname{codim}(|\beta|\setminus V_{\ell},|\beta|)>\ell$.
  • Curves in $V_{\ell}$ are either $\ell$-nodal, or reduced of geometric genus $p_{a}(\beta)-\ell$.
Note that for a curve $C\in V_{\ell}$ which is not $\ell$-nodal, its geometric genus is
(3.8) $p_{a}(\beta)-\sum_{p\in C_{\mathrm{sing}}}\delta_{p}(C)>p_{a}(\beta)-\ell.$
By (3.8) and Milnor’s formula (3.6), we have
$\sum_{p\in C_{\mathrm{sing}}}\tau_{p}(C)\leq\sum_{p\in C_{\mathrm{sing}}}\mu_{p}(C)\leq 2\sum_{p\in C_{\mathrm{sing}}}\delta_{p}(C)<2\ell.$
It remains to show Property (3). For any curve $D\in V_{\ell}$, the map $k_{\pi,D}^{\text{loc}}$ in (3.7) can be identified with
$\bar{r}_{W}:H^{0}({\mathcal{O}}_{S}(\beta))/\langle s_{D}\rangle\to H^{0}({\mathcal{O}}_{S}(\beta)|_{W})$
induced by the restriction to a 0-dimensional closed subscheme $W\subset S$ of length
$\sum_{p\in D_{\mathrm{sing}}}\tau_{p}(D)<2\ell$
by Lemma 3.3, where $s_{D}\in H^{0}({\mathcal{O}}_{S}(\beta))$ is a defining section of $D$. Hence, $\bar{r}_{W}$ is surjective by the ($2\ell-2$)-very ampleness of $\beta$. ∎
Combining the above proposition with Theorem 3.5, we can prove Theorem 1.2.
Proof of Theorem 1.2.
Let $V_{\ell}$ be as in Proposition 3.6. Take $U_{\ell}=V_{\ell}\cap|\beta|^{\mathrm{int}}$. By Property (2) of $V_{\ell}$ in Proposition 3.6, we have
$\operatorname{codim}(|\beta|^{\mathrm{int}}\setminus U_{\ell},|\beta|^{\mathrm{int}})\geq\ell+1.$
It remains to show that $h^{-1}(U_{\ell})$ is smooth. Note that $h^{-1}(U_{\ell})$ is the relative compactified Jacobian of degree $p_{a}(\beta)-1+\chi$ associated with the universal family of curves in $U_{\ell}$. This universal family is locally versal by Property (3) of $V_{\ell}$ in Proposition 3.6. Thus, the smoothness of $h^{-1}(U_{\ell})$ follows from Theorem 3.5. ∎
4. Intersection Betti numbers
In this section, we generalize the approach in [29] to prove Theorem 1.1. For any open subscheme $U\subset|\beta|$, let
$N(U)\vcentcolon=2\operatorname{codim}(|\beta|\setminus U,|\beta|)-2.$
We simply write $N(\beta)$ for $N(|\beta|^{\mathrm{int}})$.
First, we focus on a smooth open subscheme $h^{-1}(U)$ of $M_{\beta,\chi}$ and prove the following result concerning the dimensions of graded pieces of the perverse filtration induced by the restriction of $h:M_{\beta,\chi}\to|\beta|$ (see (2.4)).
Theorem 4.1.
Suppose that $|\beta|$ contains a connected smooth curve. Let $U\subset|\beta|^{\mathrm{int}}$ be an open subscheme such that $h^{-1}(U)$ is smooth, and let $h_{U}:h^{-1}(U)\to U$ be the restriction of $h$. For given $i,j\in\mathbb{Z}_{\geq 0}$, if $i+j\leq N(U)$, the divisor $\beta$ is (i1)(i-1)-very ample, and $2\dim|\beta|\geq 3i+j$, then
(4.9) $n_{h_{U}}^{i,j}=n_{\infty}^{i,j},$
where $n_{\infty}^{i,j}$ is the coefficient of $q^{i}t^{j}$ in the expansion of
(4.10) $H(q,t)\vcentcolon=(1-qt)\prod_{m\geq 1}\frac{(1+q^{m}t^{m-1})^{b_{1}(S)}(1+q^{m}t^{m+1})^{b_{1}(S)}}{(1-q^{m+1}t^{m-1})(1-q^{m}t^{m})^{b_{2}(S)}(1-q^{m-1}t^{m+1})}.$
Remark 4.2.
Recall Göttsche’s formula [8] for $\sum_{n\geq 0}\sum_{i\geq 0}b_{i}(S^{[n]})z^{i}w^{n}$:
(4.11) $G(z,w)\vcentcolon=\prod_{m\geq 1}\frac{(1+z^{2m-1}w^{m})^{b_{1}(S)}(1+z^{2m+1}w^{m})^{b_{1}(S)}}{(1-z^{2m-2}w^{m})(1-z^{2m}w^{m})^{b_{2}(S)}(1-z^{2m+2}w^{m})}.$
After the change of variables
$z=t,\quad w=\frac{q}{t},$
the product formula $H(q,t)$ in (4.10) satisfies the following equality
$\frac{H(q,t)}{1-qt}=G(z,w)\cdot\frac{1-w}{1-z^{2}}.$
By (2.4), the $k$-th Betti number of $h^{-1}(U)$ is
(4.12) $b_{k}(h^{-1}(U))=\sum_{i=0}^{k}n_{h_{U}}^{i,k-i}.$
Since the right-hand side of (1.1) is exactly $H(q,q)$, the next corollary follows immediately from Theorem 4.1 and (4.12).
Corollary 4.3.
Let $k\in\mathbb{Z}_{\geq 0}$. In the setting of Theorem 4.1, if $k\leq\min\{N(U),2\dim|\beta|/3\}$ and $\beta$ is (k1k-1)-very ample, then
$b_{k}(h^{-1}(U))=b_{k}^{\infty},$
where $b^{\infty}_{k}$ is defined by (1.1).
4.1. Strategy of the proof
The proof of Theorem 4.1 consists of the following steps.
Step 1. Relate $h^{-1}(U)$ to the relative Hilbert scheme ${\mathcal{C}}^{[n]}_{U}$. By (2.5), for any $\ell,m\in\mathbb{Z}_{\geq 0}$,
(4.13) $b_{m}({\mathcal{C}}^{[\ell]}_{U})=\sum_{i+j\leq\ell;\,i,j\geq 0}n_{h_{U}}^{i,m-i-2j}.$
Step 2. Calculate the Betti numbers of ${\mathcal{C}}^{[\ell]}_{U}$ using Proposition 2.3.
Proposition 4.4.
In the setting of Theorem 4.1, if $m\leq\min\{N(U),2\dim|\beta|-2\ell\}$ and $\beta$ is ($\ell-1$)-very ample, then
(4.14) $b_{m}({\mathcal{C}}^{[\ell]}_{U})=\sum_{n=0}^{\lfloor m/2\rfloor}b_{m-2n}(S^{[\ell]}).$
Proof.
Since $\pi^{[\ell]}:{\mathcal{C}}^{[\ell]}\to|\beta|$ has fibers of dimension $\ell$ by Proposition 2.1, we have
$\operatorname{codim}({\mathcal{C}}^{[\ell]}\setminus{\mathcal{C}}_{U}^{[\ell]},{\mathcal{C}}^{[\ell]})=\operatorname{codim}(|\beta|\setminus U,|\beta|).$
By Proposition 2.3, the $(\ell-1)$-very ampleness of $\beta$ implies that $\mathcal{C}$$^{[\ell]}$ is a projective bundle over $S^{[\ell]}$. In particular, $\mathcal{C}$$^{[\ell]}$ is irreducible and smooth. Then it follows from [29, Lemma 2.4] (or the proof of Lemma 5.6) that for $m\leq N(U)$,
$b_{m}({\mathcal{C}}^{[\ell]}_{U})=b_{m}({\mathcal{C}}^{[\ell]}).$
Combining with
$\dim{\mathcal{C}}^{[\ell]}-\dim S^{[\ell]}=\dim|\beta|-\ell\geq m/2$
by our assumption, we conclude
$b_{m}({\mathcal{C}}^{[\ell]}_{U})=b_{m}({\mathcal{C}}^{[\ell]})=b_{m}(S^{[\ell]}\times\mathbb{P}^{\dim|\beta|-\ell})=\sum_{n=0}^{\lfloor m/2\rfloor}b_{m-2n}(S^{[\ell]}),$
which proves (4.14). ∎
Since all Betti numbers of $S^{[\ell]}$ are determined by Göttsche’s formula (4.11), we can prove Theorem 4.1 by induction on $i$, combining (4.13) and (4.14).
Proof of Theorem 4.1.
We proceed by induction on $i$ to prove (4.9). For $i=0$ and $j\leq\min\{N(U),2\dim|\beta|\}$,
$n_{h_{U}}^{0,j}$ $=\dim P_{0}H^{j}(h^{-1}(U))$
[…]
which proves the induction base of (4.9). Suppose the result is true for $(i^{\prime},j^{\prime})$ with $i^{\prime}<i$ and $i^{\prime}+j^{\prime}\leq i+j$. For $m\leq i+j$, it follows by (4.13), our induction hypothesis, and (4.14) that
$n_{h_{U}}^{i,m-i}$ $=\sum_{i^{\prime}+j^{\prime}=i;\,i^{\prime},j^{\prime}\geq 0}n_{h_{U}}^{i^{\prime},m-i^{\prime}-2j^{\prime}}-\sum_{i^{\prime}<i;\,i^{\prime}+j^{\prime}=i;\,i^{\prime},j^{\prime}\geq 0}n_{h_{U}}^{i^{\prime},m-i^{\prime}-2j^{\prime}}$
$=\left(b_{m}({\mathcal{C}}^{[i]}_{U})-b_{m}({\mathcal{C}}^{[i-1]}_{U})\right)-\sum_{i^{\prime}<i;\,i^{\prime}+j^{\prime}=i;\,i^{\prime},j^{\prime}\geq 0}n_{\infty}^{i^{\prime},m-i^{\prime}-2j^{\prime}}$
$=\sum_{n=0}^{\lfloor m/2\rfloor}\left(b_{m-2n}(S^{[i]})-b_{m-2n}(S^{[i-1]})\right)-\sum_{i^{\prime}<i;\,i^{\prime}+j^{\prime}=i;\,i^{\prime},j^{\prime}\geq 0}n_{\infty}^{i^{\prime},m-i^{\prime}-2j^{\prime}}$
$=n_{\infty}^{i,m-i},$
where the last identity is essentially the same as [29, (31)] by Göttsche’s formula [8] (see also Remark 4.2). ∎
Finally, we prove our first main theorem stated in the introduction.
Proof of Theorem 1.1.
By the upper semicontinuity of fiber dimensions, we have
(4.15) $N(\beta)\geq 2c-2.$
Since $\beta$ is $\max\{2,2k-2\}$-very ample, let $U_{k}\subset|\beta|^{\mathrm{int}}$ be as in Theorem 1.2 (set $U_{k}\vcentcolon=U_{2}$ if $k\leq 1$). Using (4.15), the assumption $k\leq c-1$, and Theorem 1.2, we obtain the following properties:
  • $N(U_{k})\geq\min\{N(\beta),2k\}=2k$;
  • $h^{-1}(U_{k})$ is smooth.
By Corollary 4.3, we have
(4.16) $b_{k}(h^{-1}(U_{k}))=b_{k}^{\infty}.$
Since fibers of $h$ over integral curves have the same dimension, Theorem 1.2 implies
$\operatorname{codim}(h^{-1}(|\beta|^{\mathrm{int}})\setminus h^{-1}(U_{k}),h^{-1}(|\beta|^{\mathrm{int}}))=\operatorname{codim}(|\beta|^{\mathrm{int}}\setminus U_{k},|\beta|^{\mathrm{int}})\geq k+1.$
This, together with the assumption on $c=\operatorname{codim}(M_{\beta,\chi}\setminus h^{-1}(|\beta|^{\mathrm{int}}),M_{\beta,\chi})$, yields
$\operatorname{codim}(M_{\beta,\chi}\setminus h^{-1}(U_{k}),M_{\beta,\chi})\geq\min\{c,k+1\}>k.$
Then by [23, Theorem 6.7.4] and the smoothness of $h^{-1}(U_{k})$, we have isomorphisms
$\operatorname{IH}^{k}(M_{\beta,\chi})\cong\operatorname{IH}^{k}(h^{-1}(U_{k}))\cong H^{k}(h^{-1}(U_{k})),$
and therefore
$\dim_{\mathbb{Q}}\operatorname{IH}^{k}(M_{\beta,\chi})=b_{k}(h^{-1}(U_{k}))=b_{k}^{\infty}$
by (4.16), which completes the proof. ∎
4.2. Discussion on the general case
In general, the moduli space $M_{\beta,\chi}$ can be reducible even if $\beta$ is sufficiently positive; see [34, §5]. Moreover, even if we assume that $M_{\beta,\chi}$ is irreducible, it is usually hard to compute the codimension $c$ in Theorem 1.1. It is not necessarily true that $c$ is the same as $\operatorname{codim}(|\beta|\setminus|\beta|^{\mathrm{int}},|\beta|)$. In some nice cases where $c=\operatorname{codim}(|\beta|\setminus|\beta|^{\mathrm{int}},|\beta|)$, the asymptotic behavior of $\operatorname{IH}^{k}(M_{\beta,\chi})$ can be described more precisely, provided that $\operatorname{codim}(|\beta|\setminus|\beta|^{\mathrm{int}},|\beta|)$, or equivalently $N(\beta)$, can be calculated.
The following lemma allows us to estimate $N(\beta)$ by analyzing the decompositions of $\beta$.
Lemma 4.5.
Fix an ample divisor $\beta$ on $S$. There are finitely many pairs of classes $(\theta_{1},\theta_{2})$ in $\operatorname{Num}(S)$ such that $\theta_{i}$ ($i=1,2$) are the classes of nonzero effective divisors $C_{i}$ with
$C_{1}+C_{2}=\beta.$
Proof.
Let $D_{1},\cdots,D_{\rho}$ be an orthogonal basis for $\operatorname{Num}(S)_{\mathbb{Q}}$, where $D_{1}$ is the class of an ample divisor. By Hodge index theorem, $D_{\ell}^{2}<0$ for $1<\ell\leq\rho$. Given any $C_{1}$, $C_{2}$ as in the statement, write
$\beta\equiv\sum_{j=1}^{\rho}a_{j}D_{j},\quad C_{i}\equiv\sum_{j=1}^{\rho}a_{i,j}D_{j}$
for some $a_{j},a_{i,j}\in{\mathbb{Q}}$ ($i=1,2$), where $\equiv$ denotes the numerical equivalence. For $1<\ell\leq\rho$, let $n_{\ell}>0$ be the integer such that $A_{\ell}\vcentcolon=n_{\ell}D_{1}+D_{\ell}$ is ample. Then $D_{1}\cdot C_{i}>0$ and $A_{\ell}\cdot C_{i}>0$, which implies
$a_{i,1}>0\quad\text{and}\quad n_{\ell}a_{i,1}D_{1}^{2}+a_{i,\ell}D_{\ell}^{2}>0.$
Combined with $D_{\ell}^{2}<0$ and $C_{1}+C_{2}=\beta$, this shows that the possible values of $a_{i,j}$ ($i=1,2$; $j=1,\cdots,\rho$) are bounded since $\beta$ is fixed, and thus the result follows. ∎
By the calculations in Section 5, we ask the following question.
Question 4.6.
For any ample divisor $\beta$ on an arbitrary smooth projective surface $S$, is it true that
$\lim_{d\to+\infty}N(d\beta)=+\infty?$
5. Stabilization in the Enriques and bielliptic cases
In this section, let $S$ be an Enriques or bielliptic surface. In particular, its canonical divisor $K_{S}$ is numerically trivial. When $\beta$ is a base-point-free, ample divisor on $S$, $|\beta|$ contains a connected smooth curve by Bertini’s theorem. Denote by $|\beta|^{\mathrm{sm}}\subset|\beta|$ the open subscheme parametrizing smooth curves.
To apply Theorem 1.1, we first prove the irreducibility of the moduli spaces of one-dimensional sheaves on these surfaces.
Proposition 5.1.
Suppose that $S$ is an Enriques or bielliptic surface, and $\beta$ is a base-point-free, ample divisor satisfying $\operatorname{gcd}(\beta\cdot H,\chi)=1$. Then $M_{\beta,\chi}$ is irreducible.
Proof.
Since the open subscheme $h^{-1}(|\beta|^{\mathrm{int}})$ is irreducible and contains the smooth open subscheme $h^{-1}(|\beta|^{\mathrm{sm}})$ by [34, Proposition 2.2], there is a unique irreducible component $Y\subset M_{\beta,\chi}$ containing $h^{-1}(|\beta|^{\mathrm{int}})$. We prove the irreducibility of $M_{\beta,\chi}$ by contradiction. If $Y\not=M_{\beta,\chi}$, then take another irreducible component $Y^{\prime}\neq Y$. It follows from $Y^{\prime}\cap h^{-1}(|\beta|^{\mathrm{sm}})=\emptyset$ that $Y^{\prime}\subset h^{-1}(|\beta|\setminus|\beta|^{\mathrm{sm}})$. By [36, Corollary 1.3] and the fact that $K_{S}$ is numerically trivial, the fibers of $h:M_{\beta,\chi}\to|\beta|$ have the same dimension $p_{a}(\beta)$; therefore
(5.17) $\dim Y^{\prime}\leq\dim h^{-1}(|\beta|\setminus|\beta|^{\mathrm{sm}})<\dim|\beta|+p_{a}(\beta)=\beta^{2}+\chi({\mathcal{O}}_{S}),$
where the last equality follows from the Kodaira vanishing theorem and the Riemann–Roch formula. However, using the morphism $\operatorname{det}:M\to\operatorname{Pic}(X)$ in [11, Theorem 4.5.4]  [^2] Although [11, Theorem 4.5.4] is stated for moduli spaces of positive rank sheaves, the proof of $\dim_{[F]}M({\mathcal{Q}})\geq\dim_{[F]}M-\dim_{[\operatorname{det}F]}\operatorname{Pic}(X)$ does not rely on the positive rank assumption. and the Riemann–Roch formula, for any point ${\mathcal{F}}\in Y^{\prime}$, we have
$\dim Y^{\prime}$ $\geq\operatorname{ext}^{1}({\mathcal{F}},{\mathcal{F}})-\operatorname{ext}^{2}({\mathcal{F}},{\mathcal{F}})-h^{1}({\mathcal{O}}_{S})$
$=\beta^{2}+\chi({\mathcal{O}}_{S}),$
which contradicts (5.17). ∎
Then we assume the following property of $N(\beta)$, whose proof will be given in §5.1.2 and §5.2.2 by analyzing the Enriques and bielliptic cases separately.
Proposition 5.2.
If $\beta$ is an ample divisor on an Enriques or bielliptic surface, then
$\lim_{d\to+\infty}N(d\beta)=+\infty.$
Combining the above results with Theorem 1.1, we can prove the stabilization of intersection Betti numbers of $M_{\beta}=M_{\beta,1}$.
Theorem 5.3.
Suppose that $S$ is an Enriques or bielliptic surface. Fix $k\in\mathbb{Z}_{\geq 0}$. Given any ample divisor $\beta_{0}$ on $S$, there exists $d(\beta_{0},k)\in\mathbb{Z}_{>0}$ (depending on $\beta_{0}$ and $k$) such that
(5.18) $\dim_{\mathbb{Q}}\operatorname{IH}^{k}(M_{d\beta_{0}})=b_{k}^{\infty}\quad\text{for all integer }d\geq d(\beta_{0},k),$
where $b_{k}^{\infty}$ is defined by (1.1).
Proof.
Choose $d(\beta_{0},k)\in\mathbb{Z}_{>0}$ such that for all $d\geq d(\beta_{0},k)$, the following properties hold:
  1. (1)
    $N(d\beta_{0})\geq 2k$ (guaranteed by Proposition 5.2);
  2. (2)
    $d\beta_{0}$ is $\max\{2,2k-2\}$-very ample;
  3. (3)
    $2\dim|d\beta_{0}|\geq 3k$.
Since the fibers of $h:M_{d\beta_{0}}\to|d\beta_{0}|$ have the same dimension by [36, Corollary 1.3], we have
(5.19) $\operatorname{codim}(M_{d\beta_{0}}\setminus h^{-1}(|d\beta_{0}|^{\mathrm{int}}),M_{d\beta_{0}})$ $=\operatorname{codim}(|d\beta_{0}|\setminus|d\beta_{0}|^{\mathrm{int}},|d\beta_{0}|)$
$=1+\frac{1}{2}N(d\beta_{0})\geq k+1.$
Hence, the result follows from Proposition 5.1 and Theorem 1.1 by taking $\beta=d\beta_{0}$. ∎
Remark 5.4.
When $\beta\cdot H$ and $\chi$ are not necessarily coprime, there can be strictly semistable sheaves in $M_{\beta,\chi}$ and thus $h:M_{\beta,\chi}\to|\beta|$ may not have constant fiber dimension by [36]. In this case, one could possibly prove the irreducibility of $M_{\beta,\chi}$ by showing that the corresponding moduli stack is irreducible (cf. [19, §2.6]), but the first equality of (5.19) may fail. Therefore, to ensure the equality in (5.18) holds for $M_{d\beta_{0},\chi}$, we need an additional assumption that $d\beta_{0}\cdot H$ is coprime with $\chi$.
5.1. The case of Enriques surfaces
5.1.1. Basic properties
Recall that an Enriques surface $S$ is a minimal surface of Kodaira dimension 0 with $h^{1}({\mathcal{O}}_{S})=h^{2}({\mathcal{O}}_{S})=0$. Its second Betti number is $b_{2}(S)=10$. The following property of linear systems on $S$ is well-known (e.g., [13, Theorem]).
Lemma 5.5.
Let $D$ be a nonzero, nef, effective divisor on an Enriques surface. We have the following properties:
  1. (i)
    If $D^{2}>0$, then
    $\dim|D|=\chi({\mathcal{O}}_{S}(D))-1=\frac{D^{2}}{2}.$
  2. (ii)
    If $D^{2}=0$, then $D$ is numerically equivalent to $kE$ for some $k\in\mathbb{Z}_{>0}$ and a primitive, effective divisor $E$ with $E^{2}=0$. Moreover,
    $\dim|D|=\begin{cases}\lfloor\frac{k}{2}\rfloor&(\text{if }D=kE)\\ \lfloor\frac{k-1}{2}\rfloor&(\text{if }D=kE+K_{S}).\end{cases}$
5.1.2. Proof of Proposition 5.2 in the Enriques case
Any non-integral curves in $|d\beta|$ are in the image of $|C_{1}|\times|C_{2}|$ for finitely many decompositions $d\beta=C_{1}+C_{2}$ (up to linear or numerical equivalence) by Lemma 4.5, where $C_{1}$ and $C_{2}$ are nonzero effective divisors on $S$.
First we assume that $C_{1}$ and $C_{2}$ have nef irreducible components. To prove the lemma, we estimate all possible values of
$\dim|d\beta|-\dim|C_{1}|-\dim|C_{2}|.$
Case 1.1: $C_{1}^{2}>0$ and $C_{2}^{2}>0$. In this case, $C_{1}^{2}\geq 2$ and $C_{2}^{2}\geq 2$. By Lemma 5.5 (i) and the Hodge index theorem, we have
$\dim|d\beta|-\dim(|C_{1}|\times|C_{2}|)$ $=C_{1}\cdot C_{2}\geq\sqrt{C_{1}^{2}C_{2}^{2}}$
$\geq\sqrt{2(d^{2}\beta^{2}-2C_{1}\cdot C_{2}-2)},$
where the last inequality follows from $d^{2}\beta^{2}=C_{1}^{2}+C_{2}^{2}+2C_{1}\cdot C_{2}$. Therefore,
(5.20) $\dim|d\beta|-\dim(|C_{1}|\times|C_{2}|)=C_{1}\cdot C_{2}\geq d\sqrt{2\beta^{2}}-2.$
Case 1.2: $C_{1}^{2}=0$ and $C_{2}^{2}>0$. By Lemma 5.5 (ii), $C_{1}=sE_{1}$ or $sE_{1}+K_{S}$ for some $s\in\mathbb{Z}_{>0}$ and a primitive effective divisor $E_{1}$ with $E_{1}^{2}=0$. Moreover,
(5.21) $\dim|d\beta|-\dim(|C_{1}|\times|C_{2}|)$ $\geq C_{1}\cdot C_{2}-\lfloor\frac{s}{2}\rfloor$
$\geq s\left(d\beta\cdot E_{1}-\frac{1}{2}\right)\geq d-\frac{1}{2}.$
Case 1.3: $C_{1}^{2}=C_{2}^{2}=0$. Let $s$ and $E_{1}$ be as in Case 1.2. By Lemma 5.5 (ii), $C_{2}=tE_{2}$ or $tE_{2}+K_{S}$ for some $t\in\mathbb{Z}_{>0}$ and a primitive effective divisor $E_{2}$ with $E_{2}^{2}=0$. Then we have $C_{1}\cdot C_{2}=stE_{1}\cdot E_{2}=d^{2}\beta^{2}/2>0$ and
(5.22) $\dim|d\beta|-\dim(|C_{1}|\times|C_{2}|)$ $\geq C_{1}\cdot C_{2}-\lfloor\frac{s}{2}\rfloor-\lfloor\frac{t}{2}\rfloor\geq\frac{d^{2}\beta^{2}-(s+t)}{2}$
$\geq\frac{d^{2}\beta^{2}-(1+d^{2}\beta^{2}/2)}{2}=\frac{d^{2}\beta^{2}-2}{4}.$
It remains to consider the case where every irreducible component of $C_{1}$ (resp. $C_{2}$) is nef (resp. a (2)(-2)-curve). By Lemma 4.5, for a fixed $d$, there are finitely many such $C_{2}$. In this case, it suffices to estimate
$\dim|d\beta|-\dim|C_{1}|.$
Case 2.1: $C_{1}^{2}>0$. By Lemma 5.5 (i),
(5.23) $\dim|d\beta|-\dim|C_{1}|=\frac{d\beta\cdot C_{2}+C_{1}\cdot C_{2}}{2}\geq\frac{d}{2}.$
Case 2.2: $C_{1}^{2}=0$. Let $s$ and $E_{1}$ be as in Case 1.2. By Lemma 5.5 (ii),
(5.24) $\dim|d\beta|-\dim|C_{1}|\geq\frac{d^{2}\beta^{2}}{2}-\lfloor\frac{s}{2}\rfloor\geq\frac{s(d\beta\cdot E_{1}-1)+d\beta\cdot C_{2}}{2}\geq d-\frac{1}{2}.$
Combining (5.20)–(5.24) concludes the proof. ∎
5.1.3. Refined invariants
Assume that $S$ is a generic Enriques surface and that the class of $\beta$ in $\operatorname{Num}(S)$ is not divisible by $2$ so that the moduli space $M_{\beta}=M_{\beta,1}$ is a smooth projective Calabi–Yau variety by [30]. Denote by $n_{\beta}^{i,j}$ the number defined by (2.4) with $\phi$ taken to be $h:M_{\beta}\to|\beta|$.
In order to apply Theorem 4.1 to compute $n_{\beta}^{i,j}$, we need the following lemma which is parallel to [20, Corollary 5.3]. For the reader’s convenience, we provide a proof below.
Lemma 5.6.
Let $\beta$ be a base-point-free, ample divisor on a generic Enriques surface $S$ such that $2\nmid\beta$ in $\operatorname{Num}(S)$. When $i+j\leq N(\beta)$, we have
$n_{\beta}^{i,j}=n^{i,j}_{h^{\circ}},$
where $h^{\circ}=h_{|\beta|^{\mathrm{int}}}:h^{-1}(|\beta|^{\mathrm{int}})\to|\beta|^{\mathrm{int}}$.
Proof.
Let $W=M_{\beta}\setminus h^{-1}(|\beta|^{\mathrm{int}})$. Denote by $\iota:W\hookrightarrow M_{\beta}$ the closed embedding, and by $\jmath:h^{-1}(|\beta|^{\mathrm{int}})\hookrightarrow M_{\beta}$, $\jmath^{\circ}:|\beta|^{\mathrm{int}}\to|\beta|$ the open embeddings. Since the fibers of $h$ have the same dimension, we have
$\operatorname{codim}(W,M_{\beta})=\operatorname{codim}(|\beta|\setminus|\beta|^{\mathrm{int}},|\beta|)=1+\frac{1}{2}N(\beta).$
Consider the exact triangle
$\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}}\longrightarrow{\mathbb{Q}}_{M_{\beta}}\longrightarrow\jmath_{*}\jmath^{*}{\mathbb{Q}}_{M_{\beta}}\stackrel{{\scriptstyle+1}}{{\longrightarrow}}$
which yields another exact triangle
$h_{*}\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}}\longrightarrow h_{*}{\mathbb{Q}}_{M_{\beta}}\longrightarrow\jmath^{\circ}_{*}h^{\circ}_{*}{\mathbb{Q}}_{h^{-1}(|\beta|^{\mathrm{int}})}\stackrel{{\scriptstyle+1}}{{\longrightarrow}}.$
By the property of Verdier duals ([23, Proposition 5.3.9]) and the smoothness of $M_{\beta}$,
$\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}}\cong(D_{M_{\beta}}\iota_{*}{\mathbb{Q}}_{W})[-2\dim M_{\beta}],$
and therefore for $k\leq N(\beta)+1$,
$H^{k}(\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}})=H^{2\dim M_{\beta}-k}(W)^{\vee}=0.$
By taking the long exact sequence of cohomologies, the vanishing of $H^{k}(\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}})$ shows that for $\ell\leq N(\beta)$, the restriction map induces an isomorphism
(5.25) $H^{\ell}(M_{\beta})\cong H^{\ell}(h^{-1}(|\beta|^{\mathrm{int}})).$
Since the restriction to $h^{-1}(|\beta|^{\mathrm{int}})$ clearly preserves a relatively ample class (relative to $h$), it follows from [6, Lemma 3.3] that $n_{\beta}^{i,j}\leq n_{h^{\circ}}^{i,j}$ for $i+j\leq N(\beta)$. Then the result follows by (4.12) and (5.25). ∎
Now we can prove the stabilization of $n_{\beta}^{i,j}$ for the moduli space of one-dimensional sheaves on a generic Enriques surface.
Theorem 5.7.
Fix $i,j\in\mathbb{Z}_{\geq 0}$. Given any ample divisor $\beta_{0}$ on a generic Enriques surface $S$ such that $2\nmid\beta_{0}$ in $\operatorname{Num}(S)$, there exists $d(\beta_{0},i,j)\in\mathbb{Z}_{>0}$ (depending on $\beta_{0}$, $i$ and $j$) such that
(5.26) $n_{d\beta_{0}}^{i,j}=n_{\infty}^{i,j}\quad\text{for all integer $d\geq d(\beta_{0},i,j)$ with }2\nmid d,$
where $n_{\infty}^{i,j}$ is determined by
$\sum_{i,j\geq 0}n_{\infty}^{i,j}q^{i}t^{j}=(1-qt)\prod_{m\geq 1}\frac{1}{(1-q^{m+1}t^{m-1})(1-q^{m}t^{m})^{10}(1-q^{m-1}t^{m+1})}.$
Proof.
Choose the integer $d(\beta_{0},i,j)\in\mathbb{Z}_{>0}$ such that for all $d\geq d(\beta_{0},i,j)$, the following properties hold:
  1. (1)
    $N(d\beta_{0})\geq i+j$ (guaranteed by Proposition 5.2);
  2. (2)
    $d\beta_{0}$ is $\max\{0,i-1\}$-very ample;
  3. (3)
    $d^{2}\beta_{0}^{2}\geq 3i+j$.
Condition (3) is equivalent to $2\dim|d\beta_{0}|\geq 3i+j$ by Lemma 5.5 (i). By Lemma 5.6 and Theorem 4.1, taking $U=|d\beta_{0}|^{\mathrm{int}}$, we have
$n_{d\beta_{0}}^{i,j}=n_{h_{U}}^{i,j}=n_{\infty}^{i,j},$
which proves (5.26). ∎
Again by (4.12), the next corollary follows immediately.
Corollary 5.8.
Fix $k\in\mathbb{Z}_{\geq 0}$. Given any ample divisor $\beta_{0}$ on a generic Enriques surface $S$ such that $2\nmid\beta_{0}$ in $\operatorname{Num}(S)$, there exists $d(\beta_{0},k)\in\mathbb{Z}_{>0}$ (depending on $\beta_{0}$ and $k$) such that
$b_{k}(M_{d\beta_{0}})=b_{k}^{\infty}\quad\text{for all $d\geq d(\beta_{0},k)$ with }2\nmid d,$
where $b_{k}^{\infty}$ is determined by (1.1).
Remark 5.9.
By (5.20)–(5.22) and [35, Proposition 2.5], we can take
(5.27) $d(\beta_{0},i,j)=\max\left\{2,\,i+1,\,\left\lceil\frac{i+j+2}{2}\right\rceil,\,\left\lceil\frac{i+j+6}{2\sqrt{2\beta_{0}^{2}}}\right\rceil,\,\left\lceil\sqrt{\frac{2i+2j+6}{\beta_{0}^{2}}}\,\right\rceil\right\}$
in Theorem 1.4 (Theorem 5.7).
5.2. The case of bielliptic surfaces
5.2.1. Basic properties
Recall that a bielliptic surface $S$ is a minimal surface of Kodaira dimension 0 with $h^{1}({\mathcal{O}}_{S})=1$ and $h^{2}({\mathcal{O}}_{S})=0$. In particular, the exponential sequence reads
$0\to\operatorname{Pic}^{0}(S)\to\operatorname{Pic}(S)\to H^{2}(S,\mathbb{Z})\to 0,$
where $\operatorname{Pic}^{0}(S)=H^{1}({\mathcal{O}}_{S})/H^{1}(S,\mathbb{Z})$ is the group of algebraically trivial divisors modulo linear equivalence and has the structure of an elliptic curve. There exist two elliptic curves $A$, $B$, and an abelian group $G$ acting on $A$ and $B$, such that ([31, Proposition 1.1])
  1. (1)
    $A/G$ is elliptic and $B/G\cong\mathbb{P}^{1}$;
  2. (2)
    $S\cong(A\times B)/G$, where $G$ acts on $A\times B$ componentwise;
see [1]. We follow the convention in [31] by abuse of notation to denote by $A$ (resp. $B$) the class (in $\operatorname{Num}(S)$, $H^{2}(S,\mathbb{Z})$, or $H^{2}(S,{\mathbb{Q}})$) of a smooth fiber of the natural projection $\Psi:S\to B/G$ (resp. $\Phi:S\to A/G$). These classes satisfy
$A^{2}=B^{2}=0,\quad A\cdot B=\gamma,$
where $\gamma$ is the order of $G$. Let $\lambda A$, $\mu B$ be a basis for $\operatorname{Num}(S)$ ($\lambda,\mu\in{\mathbb{Q}}$); see [31, Theorem 1.4] for the values of $\lambda$ and $\mu$.
We need the following lemma to compute the dimension of linear systems.
Lemma 5.10 ([31, Lemma 1.3]).
Let $D$ be a divisor of numerical class $sA+tB$ with $s,t\in{\mathbb{Q}}$ on a bielliptic surface. The following properties hold:
  1. (i)
    $\chi(\mathcal{O}_{S}(D))=st\gamma$.
  2. (ii)
    $D$ is ample if and only if s>0,t>0s>0,\,t>0.
  3. (iii)
    If $D$ is ample, then $h^{0}(\mathcal{O}_{S}(D))=\chi(\mathcal{O}_{S}(D))$.
  4. (iv)
    If $H^{0}(\mathcal{O}_{S}(D))\neq 0$, then $s\geq 0,\,t\geq 0$.
5.2.2. Proof of Proposition 5.2 in the bielliptic case
The numerical class of $\beta$ in $\operatorname{Num}(S)$ is of the form $a\lambda A+b\mu B$ with $a,b\in\mathbb{Z}_{>0}$. By Lemma 5.10 (i) and (iii),
$\dim|d\beta|=\chi({\mathcal{O}}_{S}(d\beta))-1=d^{2}ab\lambda\mu\gamma-1.$
Any non-integral curve in $|d\beta|$ is in the image of a morphism from a scheme $P$ to $|d\beta|$, where $P$ parametrizes the triples $(C,D,N)$ satisfying $N\in\operatorname{Pic}^{0}(S)$, $C\in|C_{1}+N|$, and $D\in|C_{2}-N|$ for two nonzero effective divisors $C_{1}$, $C_{2}$ such that $C_{1}+C_{2}=d\beta$. By Lemma 4.5 (or Lemma 5.10), the images of finitely many such morphisms cover $|d\beta|\setminus|d\beta|^{\mathrm{int}}$. Thus, it suffices to estimate all possible values of
$\dim|d\beta|-\dim|C_{1}|\times|C_{2}|\times\operatorname{Pic}^{0}(S)=\dim|d\beta|-\dim|C_{1}|-\dim|C_{2}|-1.$
Suppose the numerical class of $C_{i}$ ($i=1,2$) is $a_{i}\lambda A+b_{i}\mu B$ ($a_{i},b_{i}\in\mathbb{Z}$). Then
$a_{1}+a_{2}=da,\quad b_{1}+b_{2}=db.$
It follows from Lemma 5.10 (iv) that $a_{i}\geq 0$ and $b_{i}\geq 0$.
Case 1: $a_{i}>0$ and $b_{i}>0$ for $i=1,2$. By Lemma 5.10 (ii) and (iii),
$\dim|C_{1}|+\dim|C_{2}|=\chi({\mathcal{O}}_{S}(C_{1}))+\chi({\mathcal{O}}_{S}(C_{2}))-2.$
Therefore, by the proof of [29, (13)] (or Case 1.1 in §5.1.2), we have
(5.28) $\dim|d\beta|-\dim|C_{1}|-\dim|C_{2}|-1=C_{1}\cdot C_{2}\geq d\sqrt{\beta^{2}}-1.$
Case 2: At least one of $a_{1},a_{2},b_{1},b_{2}$ is zero. We may assume $a_{1}=0$. Then $b_{1}>0$ since otherwise $C_{1}$ would be numerically trivial. By Lemma 5.10 (i)–(iii),
(5.29) $\dim|C_{1}|\leq\dim|A+C_{1}|=b_{1}\mu\gamma-1.$
If $b_{2}>0$, then it follows from $b_{1}>0$ and (5.29) that
(5.30) $\dim|d\beta|-\dim|C_{1}|-\dim|C_{2}|-1$ $=C_{1}\cdot C_{2}-\dim|C_{1}|$
$\geq dab_{1}\lambda\mu\gamma-(b_{1}\mu\gamma-1)$
$\geq(da\lambda-1)\mu\gamma+1\quad(\text{if }da\lambda>1).$
If $b_{2}=0$, then
$\dim|C_{2}|\leq\dim|C_{2}+B|=da\lambda\gamma-1,$
which, together with (5.29), implies
(5.31) $\dim|d\beta|-\dim|C_{1}|-\dim|C_{2}|-1$ $\geq d^{2}ab\lambda\mu\gamma-1-(db\mu\gamma-1)-(da\lambda\gamma-1)-1$
$=d^{2}ab\lambda\mu\gamma-db\mu\gamma-da\lambda\gamma.$
Combining (5.28), (5.30), and (5.31) completes the proof. ∎

### v2

@@
 Abstract.
-In this paper, we study the intersection cohomology of the moduli space of semistable one-dimensional sheaves with fixed Euler characteristic, supported in a divisor class $\beta$ on a smooth projective surface $S$.
-Assuming that this moduli space is irreducible, we prove that its intersection Betti numbers in a certain range are determined by a product formula derived from Göttsche’s formula for Betti numbers of Hilbert schemes of points on $S$.
-As an application, for Enriques and bielliptic surfaces, we show the stabilization of intersection Betti numbers of this moduli space when $\beta$ is sufficiently positive.
-In the Enriques case, we also prove a refined stabilization result concerning the perverse Hodge numbers when the moduli space is smooth.
+In this paper, we develop a unified approach to study the intersection Betti numbers of moduli spaces of one-dimensional semistable sheaves on smooth projective surfaces.
+Assuming the irreducibility of such moduli spaces, we prove that their intersection Betti numbers in a certain range of degrees coincide with the stable Betti numbers of Hilbert schemes of points.
+As an application, for minimal surfaces of Kodaira dimension 0, we show that these intersection Betti numbers stabilize in each fixed degree, which fits into the broader context of stable cohomology for moduli spaces of sheaves.
+In the case of Enriques surfaces, we also prove a refined stabilization result related to Oberdieck’s conjecture on perverse Hodge numbers.
+Key words and phrases:
+moduli spaces of sheaves, intersection cohomology, perverse filtrations, compactified Jacobians
+2020 Mathematics Subject Classification:
+14D20, 14F06 (primary), 32S60 (secondary)
 1. Introduction
-Throughout, we work over the field $\mathbb{C}$ of complex numbers. Let $S$ be a connected, smooth, projective surface with a fixed ample divisor $H$.
-In recent years, moduli spaces of one-dimensional sheaves on $S$ have attracted considerable attention.
-On the one hand, their connections with enumerative geometry and moduli spaces of Higgs bundles have led to a number of new conjectures and results; see, for instance, [3, 4, 9, 14, 15, 19, 21, 28, 29, 36].
-On the other hand, these moduli spaces exhibit not only asymptotic behavior analogous to that of moduli spaces of torsion-free sheaves, but also pathologies [34, §5] in contrast to the torsion-free case, revealing rich geometry that remains to be understood.
-Let $\chi$ be an integer and $\beta$ a nonzero effective divisor of arithmetic genus $p_{a}(\beta)$ on $S$.
-Denote by $|\beta|^{\mathrm{int}}\subset|\beta|$ the open subscheme parametrizing integral curves.
-Consider the (coarse) moduli space $M_{\beta,\chi}$ parametrizing S-equivalence classes of one-dimensional semistable (with respect to $H$) sheaves $\mathcal{F}$} on $S$ with fixed determinant $\operatorname{det}({\mathcal{F}})\cong{\mathcal{O}}_{S}(\beta)$ and Euler characteristic $\chi({\mathcal{F}})=\chi$.
-Here the stability is defined in terms of the slope
-$\mu({\mathcal{F}})=\frac{\chi({\mathcal{F}})}{c_{1}({\mathcal{F}})\cdot H}.$
-There is a morphism
-$h:M_{\beta,\chi}\to|\beta|,$
-called the Hilbert–Chow morphism, which is defined by taking the Fitting support. If $|\beta|$ contains a connected smooth curve, then the morphism $h$ is an abelian fibration.
-The moduli space $M_{\beta,\chi}$ is singular in general, for which the intersection cohomology $\operatorname{IH}^{*}(M_{\beta,\chi})$ (with $\mathbb{Q}$}-coefficients) might be a more suitable topological invariant than the ordinary cohomology.
-In this paper, we prove the following theorem about the intersection Betti numbers of $M_{\beta,\chi}$.
+In recent years, moduli spaces of one-dimensional sheaves on surfaces have attracted considerable attention.
+On the one hand, their connections with enumerative geometry and the theory of Higgs bundles have led to a number of new conjectures and results (e.g., [3, 4, 14, 20, 21, 25, 27, 35, 36, 46]).
+On the other hand, these moduli spaces exhibit not only asymptotic behavior analogous to that of moduli spaces of torsion-free sheaves, but also pathologies in contrast to the torsion-free case (see [42, §5]), revealing rich geometry that remains to be understood.
+Let $S$ be a connected, smooth, complex projective surface with a fixed ample divisor $H$.
+For a nonzero effective divisor $\beta$ on $S$ and an integer $\chi$, let $M_{\beta,\chi}$ be the coarse moduli space parametrizing S-equivalence classes of one-dimensional semistable (with respect to $H$) sheaves $\mathcal{F}$} on $S$ with fixed determinant $\det({\mathcal{F}})\cong{\mathcal{O}}_{S}(\beta)$ and Euler characteristic $\chi({\mathcal{F}})=\chi$.
+The moduli space $M_{\beta,\chi}$ admits a Hilbert–Chow morphism
+$h:M_{\beta,\chi}\to|\beta|$
+to the linear system $|\beta|$, defined by taking the Fitting support.
+Over the open subscheme $|\beta|^{\mathrm{int}}\subset|\beta|$ parametrizing integral curves, the fibers of $h$ are compactified Jacobians.
+A natural step toward understanding the geometry of $M_{\beta,\chi}$ is to determine its topological invariants.
+Since $M_{\beta,\chi}$ is singular in general, we consider its intersection cohomology, which is better-behaved than the ordinary cohomology.
+In this paper, we study the intersection Betti numbers of $M_{\beta,\chi}$, defined as the dimensions of the intersection cohomology groups $I\!H^{k}(M_{\beta,\chi})$ (with $\mathbb{Q}$}-coefficients) for $k\in\mathbb{Z}_{\geq 0}$.
+Our first main result expresses these intersection Betti numbers, in a certain range of degrees, as the coefficients $b_{k}^{\infty}$ of the formal power series
+(1.1) $\sum_{k\geq 0}b_{k}^{\infty}q^{k}=\prod_{m\geq 1}\frac{(1+q^{2m-1})^{b_{1}(S)}(1+q^{2m+1})^{b_{1}(S)}}{(1-q^{2m})^{b_{2}(S)+1}(1-q^{2m+2})},$
+where $b_{i}(S)$ denotes the $i$-th Betti number of $S$.
 Theorem 1.1.
-Fix $k\in\mathbb{Z}_{\geq 0}$ and $\chi\in\mathbb{Z}$. Suppose that $M_{\beta,\chi}$ is irreducible, and let
+Suppose that $M_{\beta,\chi}$ is irreducible, and denote by
 $c\vcentcolon=\operatorname{codim}(M_{\beta,\chi}\setminus h^{-1}(|\beta|^{\mathrm{int}}),M_{\beta,\chi})$
-denote the codimension of $M_{\beta,\chi}\setminus h^{-1}(|\beta|^{\mathrm{int}})$ in $M_{\beta,\chi}$. If the following conditions hold:
+the codimension of $M_{\beta,\chi}\setminus h^{-1}(|\beta|^{\mathrm{int}})$ in $M_{\beta,\chi}$. If
   1. (1)
-    $\beta$ is $\max\{2,2k-2\}$-very ample (see Definition 2.2);
+    $\beta$ is $\max\{1,2k-2\}$-very ample (see Definition 2.4), and
   2. (2)
-    $k\leq\min\{c-1,\frac{2}{3}\dim|\beta|\}$,
-then we have
-$\dim_{\mathbb{Q}}\operatorname{IH}^{k}(M_{\beta,\chi})=b_{k}^{\infty},$
-where $b_{k}^{\infty}$ is defined by the following identity of formal power series
-(1.1) $\sum_{k\geq 0}b_{k}^{\infty}q^{k}=\prod_{m\geq 1}\frac{(1+q^{2m-1})^{b_{1}(S)}(1+q^{2m+1})^{b_{1}(S)}}{(1-q^{2m})^{b_{2}(S)+1}(1-q^{2m+2})}.$
-By Göttsche’s formula [8], the number $b_{k}^{\infty}$ is exactly the $k$-th Betti number $b_{k}(S^{[n]})$ of the Hilbert scheme $S^{[n]}$ of $n$ points for any integer $n\geq k$.
-The connection between Betti numbers of moduli spaces of sheaves on $S$ and $b_{k}^{\infty}$ has been observed in many studies, for example, in [5, 17] for some moduli spaces of torsion-free sheaves, in [36] for moduli spaces of one-dimensional sheaves on $\mathbb{P}^{2}$, and in [29] for those of one-dimensional sheaves on any del Pezzo surface.
-Our proof of Theorem 1.1 adapts the approach in §2 and §3 of our previous joint work with Pi and Shen [29] for a more general setting. A key new ingredient in the proof is as follows.
+    $k\leq\min\left\{c-1,\frac{2}{3}\dim|\beta|\right\}$,
+then
+$\dim_{\mathbb{Q}}I\!H^{k}(M_{\beta,\chi})=b_{k}^{\infty}.$
+By Göttsche’s formula [11], the number $b_{k}^{\infty}$ is exactly the $k$-th Betti number $b_{k}(S^{[n]})$ of the Hilbert scheme $S^{[n]}$ of $n$ points on $S$ for any integer $n\geq k$.
+The connection between (intersection, virtual, or ordinary) Betti numbers of moduli spaces of sheaves on $S$ and $b_{k}^{\infty}$ has been observed in various settings.
+For some moduli spaces of torsion-free sheaves, this connection has been established, for instance, in [8, 23].
+Similar results have been obtained for moduli spaces of one-dimensional sheaves on specific surfaces such as K3 surfaces [29], the projective plane [46], and del Pezzo surfaces [36].
+The above Theorem 1.1 provides a unifying perspective on the intersection Betti numbers of moduli spaces of one-dimensional sheaves on surfaces.
+To prove Theorem 1.1, we adapt the approach in Sections 2 and 3 of our previous joint work with Pi and Shen [36] to a more general setting.
+For del Pezzo (as well as K3 and Abelian) surfaces, the open subscheme $h^{-1}(|\beta|^{\mathrm{int}})$ of $M_{\beta,\chi}$ is smooth.
+The main difficulty in generalizing the method in [36] to other cases arises from the singularities in $h^{-1}(|\beta|^{\mathrm{int}})$.
+Unlike in the torsion-free case, dimension bounds for the singular locus of $M_{\beta,\chi}$ were largely unknown, since standard techniques for torsion-free sheaves, such as those in [16, Ch.
+9], are generally inapplicable to torsion sheaves.
+We use the deformation theory of curve singularities to show that the dimension of the singular locus of $h^{-1}(|\beta|^{\mathrm{int}})$ is governed by the positivity of $\beta$, which may be of independent interest and will be essential to our proof of Theorem 1.1.
+The precise statement is as follows.
 Theorem 1.2.
-Let $\ell\geq 2$ be an integer. If $\beta$ is $(2\ell-2)$-very ample, then there exists an open subscheme $U_{\ell}\subset|\beta|^{\mathrm{int}}$ such that
+If $\beta$ is $\max\{1,2k-2\}$-very ample, then there exists an open subscheme $U_{k}\subset|\beta|^{\mathrm{int}}$ such that
   1. (1)
-    $\operatorname{codim}(|\beta|^{\mathrm{int}}\setminus U_{\ell},|\beta|^{\mathrm{int}})\geq\ell+1$;
+    $\operatorname{codim}(|\beta|^{\mathrm{int}}\setminus U_{k},|\beta|^{\mathrm{int}})\geq k+1$;
   2. (2)
-    The open subscheme $h^{-1}(U_{\ell})$ of $M_{\beta,\chi}$ is smooth.
-For simplicity of exposition, now we set $\chi=1$ and consider $M_{\beta}\vcentcolon=M_{\beta,1}$ (see Remark 5.4 for the case of arbitrary $\chi$).
-As an application of Theorem 1.1, we prove the stabilization of intersection Betti numbers of $M_{\beta}$ for Enriques and bielliptic surfaces, providing new evidence for [33, Conjecture 1.3].
-Theorem 1.3 (Theorem 5.3).
-Suppose that $S$ is an Enriques or bielliptic surface.
-Fix $k\in\mathbb{Z}_{\geq 0}$.
-Given any ample divisor $\beta_{0}$ on $S$, there exists $d(\beta_{0},k)\in\mathbb{Z}_{>0}$ (depending on $\beta_{0}$ and $k$) such that
-$\dim_{\mathbb{Q}}\operatorname{IH}^{k}(M_{d\beta_{0}})=b_{k}^{\infty}\quad\text{for all integer }d\geq d(\beta_{0},k),$
-where $b_{k}^{\infty}$ is defined by (1.1).
+    the open subscheme $h^{-1}(U_{k})$ of $M_{\beta,\chi}$ is smooth.
+Motivated by Theorem 1.1 and the results for torsion-free sheaves, we propose the following conjecture on the stabilization of intersection Betti numbers, which can be viewed as a rank 0 analogue of [8, Conjecture 1.1].
+Conjecture 1.3.
+Given any ample divisor $\beta_{0}$ on $S$ and any $k\in\mathbb{Z}_{\geq 0}$, $\chi\in\mathbb{Z}$, there exists $d(\beta_{0},k,\chi)\in\mathbb{Z}_{>0}$ such that for every integer $d\geq d(\beta_{0},k,\chi)$, the intersection Betti number $\dim_{\mathbb{Q}}I\!H^{k}(M_{d\beta_{0},\chi})$ is independent of $d$.
+If furthermore $M_{d\beta_{0},\chi}$ is irreducible, then the stable intersection Betti number is given by
+(1.2) $\dim_{\mathbb{Q}}I\!H^{k}(M_{d\beta_{0},\chi})=b_{k}^{\infty}.$
+As an application of Theorem 1.1, we prove that Conjecture 1.3 holds for all minimal surfaces of Kodaira dimension 0, namely, K3, Abelian, Enriques, and bielliptic surfaces.
+We focus primarily on the Enriques and bielliptic cases.
+Theorem 1.4.
+Suppose that $S$ is a minimal surface of Kodaira dimension 0.
+Then Conjecture 1.3 is true.
+Moreover, $d(\beta_{0},k,\chi)$ can be chosen independently of $\chi$ such that (1.2) holds for all integers $d\geq d(\beta_{0},k)\vcentcolon=d(\beta_{0},k,\chi)$.
+Remark 1.5.
+When $S$ is a K3 or an Abelian surface, the moduli space $M_{\beta,1}$ is smooth, and the statement of Conjecture 1.3 for $\chi=1$ can be deduced from known results in the literature, such as those in [45].
+In the K3 case, Mauri and Migliorini [29] proved the stabilization of intersection Betti and intersection Hodge numbers of possibly singular $M_{\beta,\chi}$.
+The Abelian case can be proved similarly using the argument in [29, §10.3].
+As in [29, 36], an analysis of $\operatorname{codim}(|d\beta_{0}|\setminus|d\beta_{0}|^{\mathrm{int}},|d\beta_{0}|)$ is needed for our proof of Theorem 1.4.
+However, our approach to determining the stable intersection Betti numbers is different from that in [29] and, in particular, does not involve the deformation equivalence of moduli spaces with Hilbert schemes of points on $S$.
+Such equivalence fails in the Enriques case (since in this case $M_{\beta,\chi}$ is odd-dimensional, e.g., [37]) and is unknown in the bielliptic case.
 Although moduli spaces of sheaves on K3 and Abelian surfaces have been extensively studied, much less is known about the cases of Enriques and bielliptic surfaces.
-In the bielliptic case, recent progress was made in [27].
-When $S$ is a generic Enriques surface and the class of $\beta$ is not divisible by $2$ in the group $\operatorname{Num}(S)$ of numerical equivalence classes of divisors on $S$, Saccà [30] proved that $M_{\beta}$ is a smooth projective Calabi–Yau variety of dimension $\beta^{2}+1$.
-In this case, although the odd cohomology of $M_{\beta}$ can be nonzero as shown in [30, Theorem 5.10], Theorem 1.3 implies that the $k$-th Betti number $b_{k}(M_{\beta})$ for fixed odd $k$ vanishes for sufficiently positive $\beta$.
-We also prove, in this setting, a refined version of Theorem 1.3 concerning the stabilization of the invariants $n_{\beta}^{i,j}$ (defined by (2.4) with $\phi$ taken to be $h:M_{\beta}\to|\beta|$) which are conjectured to refine the Gromov–Witten/Pandharipande–Thomas invariants of the local surface $\operatorname{Tot}(K_{S})$ [10, 12, 21] and are the shifted perverse Hodge numbers considered in [28].
-Theorem 1.4 (Theorem 5.7).
+In the bielliptic case, some progress was made in [34].
+When $S$ is a generic Enriques surface, assuming that $\beta^{2}>0$ and the class of $\beta$ is not divisible by $2$ in the group $\operatorname{Num}(S)$ of numerical equivalence classes of divisors on $S$, Saccà [37] proved that $M_{\beta}\vcentcolon=M_{\beta,1}$ is a smooth projective Calabi–Yau variety of dimension $\beta^{2}+1$ and computed its first few Betti numbers.
+In this case, although the odd cohomology of $M_{\beta}$ can be nonzero as shown in [37, Theorem 5.10], our Theorem 1.4 implies that for fixed odd $k$, the $k$-th Betti number $b_{k}(M_{\beta})$ vanishes if $\beta$ is sufficiently positive.
+We also prove, in this setting, a refined version (Theorem 1.6 below) of Theorem 1.4 on the stabilization of the invariants $n_{\beta}^{i,j}$ (defined by (2.6) with $\varphi$ taken to be $h:M_{\beta}\to|\beta|$).
+These invariants are conjectured to refine the Gromov–Witten/Pandharipande–Thomas invariants of the local surface $\operatorname{Tot}(K_{S})$ [15, 18, 27] and are the shifted perverse Hodge numbers considered in [35].
+Theorem 1.6.
 Suppose that $S$ is a generic Enriques surface.
-Fix $i,j\in\mathbb{Z}_{\geq 0}$.
-Given any ample divisor $\beta_{0}$ on $S$ such that $2\nmid\beta_{0}$ in $\operatorname{Num}(S)$, there exists $d(\beta_{0},i,j)\in\mathbb{Z}_{>0}$ (depending on $\beta_{0}$, $i$ and $j$) such that
-$n_{d\beta_{0}}^{i,j}=n_{\infty}^{i,j}\quad\text{for all integer $d\geq d(\beta_{0},i,j)$ with }2\nmid d,$
+Given any ample divisor $\beta_{0}$ on $S$ with $2\nmid\beta_{0}$ in $\operatorname{Num}(S)$ and any $i,j\in\mathbb{Z}_{\geq 0}$, there exists $d(\beta_{0},i,j)\in\mathbb{Z}_{>0}$ such that for all odd integers $d\geq d(\beta_{0},i,j)$,
+$n_{d\beta_{0}}^{i,j}=n_{\infty}^{i,j},$
 where $n_{\infty}^{i,j}$ is defined by the following identity of formal power series
@@
 Remark 1.7.
-Under the assumption of [28, Conjecture A], Oberdieck proved a stronger stabilization result [28, Proposition 1.1] for the shifted perverse Hodge numbers.
-We note that the product formula (1.2) coincides with the one in [28, Proposition 1.1].
-It remains open whether our approach, which does not rely on [28, Conjecture A], can be modified to show $n_{\beta}^{i,j}=n^{i,j}_{\infty}$ if $i<\beta^{2}/4$ and $j<\beta^{2}/4-1$.
-An explicit expression for $d(\beta_{0},i,j)$ in Theorem 1.4 is given in (5.27).
+Under the assumption of [35, Conjecture A], Oberdieck proved a stronger stabilization result [35, Proposition 1.1] for the shifted perverse Hodge numbers.
+We note that the product formula (1.3) coincides with the one in [35, Proposition 1.1].
+It remains open whether our approach, which does not rely on [35, Conjecture A], can be modified to show $n_{\beta}^{i,j}=n^{i,j}_{\infty}$ if $i<\beta^{2}/4$ and $j<\beta^{2}/4-1$.
+An explicit expression for $d(\beta_{0},i,j)$ in Theorem 1.6 will be given in (5.24).
 The rest of this paper is organized as follows.
-In Section 2, we collect some notation, definitions, and facts that will be useful later.
-Section 3 is devoted to the local versality of the universal family of curves in high codimension and thus to a proof of Theorem 1.2.
+In Section 2, we collect some notation, definitions, and results needed later.
+Section 3 is devoted to the local versality of the universal family of curves in high codimension and contains the proof of Theorem 1.2.
 In Section 4, we prove Theorem 1.1 and discuss some difficulties arising in the general case.
-Finally, in Section 5, we study in detail the cases of Enriques and bielliptic surfaces and prove Theorems 1.3 and 1.4.
+Finally, in Section 5, we study in detail the cases of minimal surfaces of Kodaira dimension 0 to prove Theorems 1.4 and 1.6, concluding with a discussion of related open questions.
 Conventions
-A variety is an integral and separated scheme of finite type.
+All schemes are assumed to be of finite type over the field $\mathbb{C}$ of complex numbers.
+A variety is an integral and separated scheme.
 A curve is a projective scheme of pure dimension $1$.
+A surface is a variety of dimension $2$.
 By a point of a scheme, we always mean a closed point.
@@
 $i$-th Ext group $\operatorname{Ext}^{i}({\mathcal{E}},{\mathcal{E}})$) as a $\mathbb{C}$-vector space.
-Acknowledgments
-The authors thank Weite Pi and Junliang Shen for helpful discussions.
-The second author is grateful to her advisor Jun Li for his constant support, to János Kollár and Claire Voisin for inspiring discussions on the deformation theory of singularities of planar curves, and to Georg Oberdieck for suggesting considering bielliptic surfaces.
-The first author is partially supported by the Fundamental Research Funds for the Central Universities.
-The second author is supported by the NSFC grants (No.
-12121001 and No.
-12425105) and LMNS.
 2. Preliminaries
-This section gathers the notation and background material that will be essential later.
-2.1. Relative Hilbert schemes of points
-Let $\pi:{\mathcal{C}}\to|\beta|$ be the universal family of curves in $|\beta|$. For $n\in\mathbb{Z}_{\geq 0}$, denote by
-$\pi^{[n]}:{\mathcal{C}}^{[n]}\to|\beta|$
-the relative Hilbert scheme of $n$ points on the fibers of $\pi$.
-For $C\in|\beta|$, the fiber of $\pi^{[n]}$ over $C$ is the Hilbert scheme $C^{[n]}$ parametrizing 0-dimensional, length $n$ closed subschemes of $C$.
-Note that $\pi^{[0]}:|\beta|\to|\beta|$ is the identity and $\pi=\pi^{[1]}:{\mathcal{C}}\to|\beta|$.
-For any open subscheme $U\subset|\beta|$, write ${\mathcal{C}}_{U}^{[n]}$ for $(\pi^{[n]})^{-1}(U)$.
-The following property of $\pi^{[n]}$ is a corollary of [18, Theorem 1.1] (see also [34, Proposition 2.7]).
+2.1. Moduli spaces of one-dimensional sheaves
+We first review some basic notions concerning the moduli spaces studied in this paper and then discuss some properties regarding the Hilbert–Chow morphism and irreducibility.
+Fix a smooth polarized surface $(S,H)$ as in Section 1.
+Recall that the dimension of a sheaf is the dimension of its support.
+A one-dimensional sheaf $\mathcal{F}$} on $S$ is called semistable (resp.
+stable) (with respect to $H$) if every nonzero subsheaf $\mathcal{G}$$\subsetneq{\mathcal{F}}$ is one-dimensional with
+$\frac{\chi({\mathcal{G}})}{c_{1}({\mathcal{G}})\cdot H}\leq(\text{resp.}<)\,\frac{\chi({\mathcal{F}})}{c_{1}({\mathcal{F}})\cdot H}.$
+Two semistable sheaves are called S-equivalent if their Jordan–Hölder filtrations have isomorphic factors up to reordering.
+Each S-equivalence class is represented by a unique polystable sheaf, i.e., a semistable sheaf which is the direct sum of stable sheaves.
+In general, the coarse moduli space $M_{\beta,\chi}$ considered in this paper depends on the ample divisor $H$, but we have omitted $H$ from the notation since $H$ is fixed throughout.
+Let ${\mathcal{M}}_{\beta,\chi}$ be the moduli stack of one-dimensional semistable sheaves on $S$ with $\det({\mathcal{F}})\cong{\mathcal{O}}_{S}(\beta)$ and $\chi({\mathcal{F}})=\chi$.
+Choose a large integer $n$ such that for any semistable sheaf $\mathcal{F}$} in ${\mathcal{M}}_{\beta,\chi}$, the sheaf ${\mathcal{F}}(nH)$ is globally generated and $h^{j}({\mathcal{F}}(nH))=0$ for $j>0$.
+Set $N\vcentcolon=(\beta\cdot H)n+\chi$.
+Denote by $Q$ the Quot scheme parametrizing quotients ${\mathcal{O}}_{S}(-nH)^{\oplus N}\to{\mathcal{E}}$, where $\mathcal{E}$} is semistable with fixed Hilbert polynomial $\chi({\mathcal{E}}(mH))=(\beta\cdot H)m+\chi$ and the induced map $H^{0}(S,{\mathcal{O}}_{S}^{\oplus N})\to H^{0}(S,{\mathcal{E}}(nH))$ is an isomorphism.
+Let $\operatorname{Pic}(S)$ denote the Picard scheme of $S$. There is a morphism
+$\det:Q\to\operatorname{Pic}(S)$
+sending ${\mathcal{O}}_{S}(-nH)^{\oplus N}\to{\mathcal{E}}]$ to $\det({\mathcal{E}})$. Then ${\mathcal{M}}_{\beta,\chi}$ has a quotient stack presentation
+${\mathcal{M}}_{\beta,\chi}\simeq[Q_{\beta}/\operatorname{GL}(N)],$
+where $Q_{\beta}\vcentcolon=\det^{-1}({\mathcal{O}}_{S}(\beta))$ and $\operatorname{GL}(N)$ is the general linear group acting naturally on $Q_{\beta}$.
+The natural morphism
+$\nu:{\mathcal{M}}_{\beta,\chi}\to M_{\beta,\chi}$
+is the good moduli space for ${\mathcal{M}}_{\beta,\chi}$. Denote by
+$\tilde{h}=h\circ\nu:{\mathcal{M}}_{\beta,\chi}\to|\beta|$
+the stacky Hilbert–Chow morphism.
+Since the fibers of $h$ over the locus $|\beta|^{\mathrm{int}}$ of integral curves are compactified Jacobians and have dimension $p_{a}(\beta)$, where $p_{a}(\beta)$ is the arithmetic genus of $\beta$, the fibers of $\tilde{h}$ over $|\beta|^{\mathrm{int}}$ have dimension $p_{a}(\beta)-1$.
+Whether the fiber dimension of $\tilde{h}$ jumps depends on numerical properties of the canonical divisor $K_{S}$.
+Lemma 2.1 ([46, Corollary 1.3]).
+If $K_{S}$ is nef and $|\beta|^{\mathrm{int}}\neq\varnothing$, then for any curve $C\in|\beta|$, the fiber $\tilde{h}^{-1}(C)$ has dimension $p_{a}(\beta)-1$.
+This property of $\tilde{h}$ gives constraints on the geometry of $M_{\beta,\chi}$. The next proposition generalizes [25, Theorem 2.3] and [46, Theorem 1.5] beyond the del Pezzo case.
 Proposition 2.2.
-The morphism $\pi^{[n]}:{\mathcal{C}}^{[n]}\to|\beta|$ has fibers of the same dimension $n$.
-There is a natural morphism $\sigma_{n}:{\mathcal{C}}^{[n+1]}\to S^{[n+1]}$ defined by
-(2.3) $Z\subset C]\mapsto[Z\subset S],$
-where $C\in|\beta|$ and $Z\subset C$ is a 0-dimensional closed subscheme of length $n+1$.
-The fiber of $\sigma_{n}$ over $Z\subset S]\in S^{[n+1]}$ is the projective space $\mathbb{P}(H^{0}(S,{\mathcal{I}}_{Z}\otimes{\mathcal{O}}_{S}(\beta))^{\vee})$, where ${\mathcal{I}}_{Z}$ is the ideal sheaf of $Z\subset S$.
-2.2. Positivity conditions
-We recall the following notion of $k$-very ampleness, which generalizes base-point-freeness (0-very ampleness) and very ampleness ($1$-very ampleness).
+Suppose that $K_{S}$ is nef and $|\beta|$ contains a smooth connected curve. If $\dim|\beta|=\chi({\mathcal{O}}_{S}(\beta))-1$, then
+  1. (i)
+    the moduli space $M_{\beta,\chi}$ is irreducible;
+  2. (ii)
+    all fibers of $h:M_{\beta,\chi}\to|\beta|$ have dimension $p_{a}(\beta)$.
+Proof.
+(i) It suffices to show that $Q_{\beta}$ is irreducible.
+We use the deformation theory of sheaves and quotients (see [9, Theorem 6.4.9, Proposition 6.5.1] and [16, 2.A.8, Theorem 4.5.3]).
+For any $\xi\in Q_{\beta}$ representing the short exact sequence
+(2.4) $0\to{\mathcal{K}}\to{\mathcal{O}}_{S}(-nH)^{\oplus N}\to{\mathcal{F}}\to 0,$
+the dimension of $Q_{\beta}$ at $\xi$, denoted by $\dim_{\xi}Q_{\beta}$, satisfies (cf. [16, Proposition 4.5.9])
+(2.5) $\dim_{\xi}Q_{\beta}$ $\geq\dim_{\xi}Q-\dim\operatorname{Pic}(S)$
+$\geq\operatorname{ext}^{0}({\mathcal{K}},{\mathcal{F}})-(\operatorname{ext}^{2}({\mathcal{F}},{\mathcal{F}})-h^{2}({\mathcal{O}}_{S}))-h^{1}({\mathcal{O}}_{S})$
+$=N^{2}+\beta^{2}+\chi({\mathcal{O}}_{S})-1,$
+where the last equality follows from the long exact sequence of Ext groups for (2.4) and the Riemann–Roch formula.
+By [42, Proposition 2.2], the open subscheme $h^{-1}(|\beta|^{\mathrm{int}})$ is irreducible and contains the smooth open subscheme $h^{-1}(|\beta|^{\mathrm{sm}})$, where $|\beta|^{\mathrm{sm}}\subset|\beta|$ is the open subscheme parametrizing smooth curves.
+Then by [16, Corollary 4.3.5], the open subscheme $Q_{\beta}^{\mathrm{int}}\subset Q_{\beta}$ parametrizing those quotient sheaves with integral support is irreducible and contains the smooth open subscheme $Q_{\beta}^{\mathrm{sm}}$ parametrizing those quotient sheaves with smooth support.
+There is a unique irreducible component $Y\subset Q_{\beta}$ containing $Q_{\beta}^{\mathrm{int}}$.
+If $Y\not=Q_{\beta}$, by taking an irreducible component $Y^{\prime}\neq Y$ and noting that $Y^{\prime}\cap Q_{\beta}^{\mathrm{sm}}=\varnothing$, we deduce from Lemma 2.1 and [43, Tag 0DS4] that
+$\dim Y^{\prime}<\dim\operatorname{GL}(N)+\dim|\beta|+(p_{a}(\beta)-1)=N^{2}+\beta^{2}+\chi({\mathcal{O}}_{S})-1,$
+which contradicts (2.5). Hence, $Q_{\beta}=Y$ is irreducible, and so is $M_{\beta,\chi}$.
+(ii) By the upper semicontinuity of the fiber dimension of $h$, $\dim h^{-1}(C)\geq p_{a}(\beta)$ for any curve $C\in|\beta|$.
+Note that ${\mathcal{M}}_{\beta,\chi}$ is also irreducible since $Q_{\beta}$ is.
+Consider the morphism $\nu:{\mathcal{M}}_{\beta,\chi}\to M_{\beta,\chi}$ and its restriction $\nu|_{\tilde{h}^{-1}(C)}:\tilde{h}^{-1}(C)\to h^{-1}(C)$.
+By Lemma 2.1 and [43, Tag 0DS4], we have
+$\dim\tilde{h}^{-1}(C)-\dim h^{-1}(C)=(p_{a}(\beta)-1)-\dim h^{-1}(C)\geq-1=\dim{\mathcal{M}}_{\beta,\chi}-\dim M_{\beta,\chi},$
+which forces $\dim h^{-1}(C)=p_{a}(\beta)$. ∎
+Remark 2.3.
+If $\beta$ is an ample divisor on a K3 (resp.
+an Abelian) surface $S$ satisfying $\operatorname{codim}(|\beta|\setminus|\beta|^{\mathrm{int}},|\beta|)\geq 2$, Proposition 2.2 (ii) can also be obtained from the symplectic geometry of $M_{\beta,\chi}$: [^1] The authors are grateful to Mirko Mauri for pointing out this alternative approach and the relevant references via private communication.
+Using [5, Propositions 3.5, 3.14], one can show that $M_{\beta,\chi}$ (resp.
+$K_{\beta,\chi}$) is a primitive symplectic variety (e.g., [5, Definition 3.3]) admitting a Lagrangian fibration over $|\beta|$ (with no restrictions on $H$ or $\chi$), and thus all fibers of $h:M_{\beta,\chi}\to|\beta|$ have dimension $p_{a}(\beta)$ by [38, Theorem 3].
+Here, when $S$ is an Abelian surface with identity $0_{S}$, we denote by $K_{\beta,\chi}$ the fiber $\widehat{\varphi}^{\,-1}(0_{S})$ of the isotrivial fibration $\widehat{\varphi}:M_{\beta,\chi}\to S$ (cf.
+[45, Definition 4.1]).
+2.2. Intersection cohomology and perverse filtrations
+For a quasi-projective variety $X$, let $\operatorname{IC}_{X}$ denote the intersection cohomology complex of $X$ with $\mathbb{Q}$}-coefficients and middle perversity.
+The intersection cohomology groups of $X$ are defined as the hypercohomology groups of $\operatorname{IC}_{X}$:
+$I\!H^{k}(X)\vcentcolon={\mathbb{H}}^{k-\dim X}(X,\operatorname{IC}_{X}).$
+Note that $I\!H^{k}(X)$ can also be defined if $X$ is allowed to be non-reduced (by the topological invariance of intersection cohomology) or reducible (even not pure dimensional, as shown in [6, §4.6]).
+When $X$ is smooth, $I\!H^{k}(X)$ can be identified with the rational cohomology group $H^{k}(X)$.
+Let $\varphi:X\to Y$ be a proper morphism between quasi-projective varieties. We assume that $\dim X=a$, $\dim Y=b$, and all fibers of $\varphi$ have dimension aba-b. The perverse filtration
+$P_{0}I\!H^{k}(X)\subset P_{1}I\!H^{k}(X)\subset\dots\subset I\!H^{k}(X)$
+induced by $\varphi$ is an increasing filtration on $I\!H^{k}(X)$, governed by the topology of the morphism $\varphi$. It is defined by
+$P_{i}I\!H^{k}(X)\vcentcolon=\mathrm{Im}\left\{{\mathbb{H}}^{k-b}(Y,{{}^{\mathbf{p}}\tau_{\leq i}}(R\varphi_{*}\operatorname{IC}_{X}[b-a]))\to{\mathbb{H}}^{k-b}(Y,R\varphi_{*}\operatorname{IC}_{X}[b-a])\right\},$
+where $^{\mathbf{p}}\tau_{\leq\bullet}$ is the perverse truncation functor.
+We denote the dimension of the graded piece $\mathrm{Gr}_{i}^{P}I\!H^{i+j}(X)\vcentcolon=P_{i}I\!H^{i+j}(X)/P_{i-1}I\!H^{i+j}(X)$ by
+(2.6) $n_{\varphi}^{i,j}\vcentcolon=\dim_{\mathbb{Q}}\mathrm{Gr}_{i}^{P}I\!H^{i+j}(X).$
+When $X$ is smooth, the decomposition theorem [2] applied to $\varphi:X\to Y$ yields
+$R\varphi_{*}{\mathbb{Q}}_{X}[b]\simeq\bigoplus_{i=0}^{2a-2b}\mathcal{P}_{i}[-i],$
+where $\mathcal{P}_{i}$ are semisimple perverse sheaves on $Y$; in this case, $n_{\varphi}^{i,j}=\dim_{\mathbb{Q}}{\mathbb{H}}^{j-b}(Y,{\mathcal{P}}_{i})$.
+For our purposes, we mainly consider the Hilbert–Chow morphism $h:M_{\beta,\chi}\to|\beta|$ and its restrictions.
+The invariants $n_{\varphi}^{i,j}$ with $\varphi$ taken to be (the restrictions of) $h$ will play a central role in Section 4.
+2.3. Positivity conditions for divisors
+To study the asymptotic behavior of $M_{\beta,\chi}$, we recall the definition of $k$-very ampleness for divisors on $S$, which generalizes base-point-freeness (0-very ampleness) and very ampleness ($1$-very ampleness).
 Definition 2.4.
-An effective divisor $\beta$ on $S$ is called $k$-very ample $(k\in\mathbb{Z}_{\geq 0})$ if for any 0-dimensional closed subscheme $Z\subset S$ of length $k+1$, the restriction map
+An effective divisor $\beta$ on $S$ is called $k$-very ample if for any 0-dimensional closed subscheme $Z\subset S$ of length $k+1$, the restriction map
 $r_{Z}:H^{0}(S,{\mathcal{O}}_{S}(\beta))\to H^{0}(Z,{\mathcal{O}}_{S}(\beta)|_{Z})$
 is surjective. We adopt the convention that every divisor is (1)(-1)-very ample.
-The definition immediately implies two properties, which will be used in the sequel:
+When $h^{1}({\mathcal{O}}_{S}(\beta))=0$, the restriction map $r_{Z}$ is surjective if and only if $h^{1}({\mathcal{I}}_{Z}\otimes{\mathcal{O}}_{S}(\beta))=0$ by the long exact sequence of cohomology groups for the short exact sequence
+$0\to{\mathcal{I}}_{Z}\otimes{\mathcal{O}}_{S}(\beta)\to{\mathcal{O}}_{S}(\beta)\to{\mathcal{O}}_{S}(\beta)|_{Z}\to 0,$
+where ${\mathcal{I}}_{Z}$ is the ideal sheaf of $Z\subset S$. Definition 2.4 immediately implies two properties, which will be used in the subsequent sections:
   1. (1)
@@
   2. (2)
-    If $\beta$ is ample, then for any $k\in\mathbb{Z}_{\geq 0}$, there exists $d_{k}\in\mathbb{Z}_{>0}$ such that the multiple $d\beta$ is $k$-very ample for every integer $d\geq d_{k}$.
-A well-known connection between $k$-very ampleness and the structure of ${\mathcal{C}}^{[k+1]}$ as a scheme over $S^{[k+1]}$ via $\sigma_{k}$ is as follows (e.g., [34, Proposition 2.6]).
-Proposition 2.3.
-If $\beta$ is $k$-very ample, then ${\mathcal{C}}^{[k+1]}$ is a projective bundle over $S^{[k+1]}$.
-2.3. Perverse filtrations
-Next we recall some generalities of perverse filtrations.
-Let $\phi:X\to Y$ be a proper morphism between smooth quasiprojective varieties.
-We assume that $\dim X=a$, $\dim Y=b$, and all fibers of $\phi$ have dimension aba-b.
-The perverse filtration
-$P_{0}H^{m}(X)\subset P_{1}H^{m}(X)\subset\dots\subset H^{m}(X)$
-induced by $\phi$ is an increasing filtration on the rational cohomology $H^{*}(X)$, governed by the topology of the morphism $\phi$. It is defined to be
-$P_{k}H^{m}(X)\vcentcolon=\mathrm{Im}\left\{H^{m-b}(Y,{{}^{\mathbf{p}}\tau_{\leq k}}(R\phi_{*}{\mathbb{Q}}_{X}[b]))\to H^{m-b}(Y,R\phi_{*}{\mathbb{Q}}_{X}[b])\right\}\subset H^{m}(X),$
-where $^{\mathbf{p}}\tau_{\leq\bullet}$ is the perverse truncation functor [2]. If we apply the decomposition theorem to $\phi:X\to Y$, we obtain
-$R\phi_{*}{\mathbb{Q}}_{X}[b]\simeq\bigoplus_{i=0}^{2a-2b}\mathcal{P}_{i}[-i]\in D^{b}_{c}(Y)$
-with $\mathcal{P}_{i}$ a semisimple perverse sheaf on $Y$. There is an identity
-$P_{k}H^{m}(X)=\mathrm{Im}\Big\{H^{m-b}(Y,\bigoplus_{i=0}^{k}\mathcal{P}_{i}[-i])\to H^{m}(X)\Big\}.$
-The dimension of the graded piece $\mathrm{Gr}_{i}^{P}H^{i+j}(X)\vcentcolon=P_{i}H^{i+j}(X)/P_{i-1}H^{i+j}(X)$ is denoted by
-(2.4) $n_{\phi}^{i,j}\vcentcolon=\dim_{\mathbb{Q}}\mathrm{Gr}_{i}^{P}H^{i+j}(X)=\dim_{\mathbb{Q}}H^{j-b}(Y,{\mathcal{P}}_{i}).$
-Note that for the Hilbert–Chow morphism $h:M_{\beta,\chi}\to|\beta|$, the fibers of $h$ over integral curves in $|\beta|$ are the compactified Jacobians of degree $p_{a}(\beta)-1+\chi$ and have the same dimension $p_{a}(\beta)$.
-Later, we will use the following result of Maulik–Yun [22, Theorem 2.13] and Migliorini–Shende [24, Corollary 2].
-Theorem 2.4 ([22, 24] [^1] The result of [22, 24] can be applied here since the smoothness of $h^{-1}(U)$ implies that all relative compactified Jacobians $\overline{J}_{U}^{e}$ ($e\in\mathbb{Z}$) are smooth by the existence of a section of the universal curve (étale locally on $U$, avoiding the singularities of fibers) and so are all relative Hilbert schemes ${\mathcal{C}}^{[\ell]}_{U}$ ($\ell\in\mathbb{Z}_{\geq 0}$) by [32, Proposition 14].).
-Suppose that $|\beta|$ contains a connected smooth curve.
-Let $U\subset|\beta|^{\mathrm{int}}$ be an open subscheme such that $h^{-1}(U)$ is smooth, and let $P_{\bullet}H^{*}(h^{-1}(U))$ be the perverse filtration induced by the restriction $h_{U}:h^{-1}(U)\to U$ of $h$.
-For any $\ell,m\in\mathbb{Z}_{\geq 0}$, we have an isomorphism of vector spaces
-(2.5) $H^{m}\left({\mathcal{C}}^{[\ell]}_{U}\right)\cong\sum_{i+j\leq\ell;\,i,j\geq 0}\mathrm{Gr}_{i}^{P}H^{m-2j}\left(h^{-1}(U)\right).$
-Remark 2.5.
-When $\ell=0$, the isomorphism (2.5) reads
-$H^{m}(U)\cong\mathrm{Gr}_{0}^{P}H^{m}\left(h^{-1}(U)\right).$
-This isomorphism can be obtained as a consequence of Ngô’s support theorem [26, Théorème 7.2.1].
-This theorem shows that $h_{U}:h^{-1}(U)\to U$ has full support; see [22, Theorem 2.4, §2.5] or [24, Corollary 9].
-More precisely, there is an isomorphism
-$Rh_{U*}{\mathbb{Q}}_{h^{-1}(U)}[\dim U]\cong\bigoplus_{i=0}^{2p_{a}(\beta)}\operatorname{IC}(\wedge^{i}R^{1}\pi^{s}_{*}{\mathbb{Q}}_{{\mathcal{C}}^{\mathrm{sm}}})[-i],$
-where $\pi^{s}:{\mathcal{C}}^{\mathrm{sm}}\to|\beta|^{\mathrm{sm}}$ is the universal family of smooth curves in $|\beta|$.
+    If $\beta$ is ample, then for any $k\in\mathbb{Z}_{\geq 0}$, there exists $d_{k}\in\mathbb{Z}_{>0}$ such that $d\beta$ is $k$-very ample for every integer $d\geq d_{k}$.
 3. Singular locus of relative compactified Jacobians
-The goal of this section is to prove Theorem 1.2 about the smoothness of a big open subscheme of $h^{-1}(|\beta|^{\mathrm{int}})$.
-We begin by reviewing some basic notions of plane curve singularities and deformation theory.
-3.1. Singularities and deformations of planar curves
+The goal of this section is to prove Theorem 1.2. We begin by reviewing some basic notions of plane curve singularities and deformation theory.
+3.1. Singularities and deformations of locally planar curves
 3.1.1. Singularity invariants
-We first recall some basic invariants of isolated singularities in planar curves.
+We first recall some analytic invariants of isolated singularities in locally planar curves.
 Definition 3.1.
 Let $f\in\mathbb{C}[[x,y]]$ be a convergent power series.
-  1. (1)
+  1. (i)
     The numbers μ(f):=dim_CC[[x,y]]/(∂_x f,∂_y f), τ(f):=dim_CC[[x,y]]/(f,∂_x f,∂_y f) are called the Milnor number and Tjurina number of $f$, respectively.
-  2. (2)
+  2. (ii)
     If $f$ is reduced, let O=C[[x,y]]/(f)↪O be the normalization of $\mathcal{O}$}. Then the number δ(f):=dim_CO/O is called the $\delta$-invariant of $f$.
 The Milnor number and the $\delta$-invariant are related by the following formula of Milnor.
-Proposition 3.2 ([25, §10]).
+Proposition 3.2 ([32, §10]).
 Let $f\in(x,y)\subset\mathbb{C}[[x,y]]$ be a reduced, convergent power series. Denote by $r(f)$ the number of irreducible factors of $f$. Then
@@
 For any point $p\in C_{\mathrm{sing}}$, there exists a polynomial $f\in(x,y)\subset\mathbb{C}[x,y]$ without multiple factors such that the singularity of $C$ at $p$ is analytically isomorphic to that of $\{f=0\}$ at $(0,0)$, i.e.,
-$\hat{\mathcal{O}}_{C,p}\cong\mathbb{C}[[x,y]]/(f).$
-Thus, the Milnor and Tjurina numbers, and the $\delta$-invariant of $C$ at $p$ can be defined as those of $f$ in Definition 3.1, denoted by $\mu_{p}(C)$, $\tau_{p}(C)$, and $\delta_{p}(C)$, respectively.
-For completeness, we give a proof of the following standard interpretation of the total Tjurina number of $C$ in terms of the Ext sheaf ${\mathcal{E}xt}^{1}(\Omega_{C},{\mathcal{O}}_{C})$, where $\Omega_{C}$ is the cotangent sheaf of $C$.
+$\widehat{\mathcal{O}}_{C,p}\cong\mathbb{C}[[x,y]]/(f).$
+Thus, the Milnor number, the Tjurina number, and the $\delta$-invariant of $C$ at $p$ can be defined as those of $f$ in Definition 3.1, denoted by $\mu_{p}(C)$, $\tau_{p}(C)$, and $\delta_{p}(C)$, respectively.
+For completeness, we give a proof of the following standard interpretation (cf.
+[13, Ch.
+II, Exercise 1.4.2]) of the total Tjurina number of $C$ in terms of the Ext sheaf ${\mathcal{E}xt}^{1}(\Omega_{C},{\mathcal{O}}_{C})$, where $\Omega_{C}$ is the cotangent sheaf of $C$.
 Lemma 3.3.
@@
 The lemma can be proved locally around $p$.
-For $R\vcentcolon=\mathbb{C}[x,y]/(f)$, we have $\Omega_{R}=(Rdx\oplus Rdy)/(\partial_{x}fdx+\partial_{y}fdy)$.
-By the conormal exact sequence
+For $R\vcentcolon=\mathbb{C}[x,y]/(f)$, we have $\Omega_{R}\cong(Rdx\oplus Rdy)/(\partial_{x}fdx+\partial_{y}fdy)$.
+Using the conormal exact sequence
 $0\to(f)/(f^{2})\to R^{\oplus 2}\to\Omega_{R}\to 0,$
-we conclude that $\operatorname{Ext}^{1}(\Omega_{R},R)=R/(\partial_{x}f,\partial_{y}f)$, from which the result follows. ∎
+we obtain $\operatorname{Ext}^{1}(\Omega_{R},R)\cong R/(\partial_{x}f,\partial_{y}f)$, from which the result follows. ∎
 3.1.2. Smoothness from local versality
-Let $\pi_{B}:{\mathcal{C}}_{B}\to B$ be a family of reduced planar curves, i.e., a proper and flat morphism whose fibers are reduced curves with at worst planar singularities.
+Let $\pi_{B}:{\mathcal{C}}_{B}\to B$ be a family of reduced, locally planar curves, i.e., a proper and flat morphism whose fibers are reduced curves with at worst planar singularities.
 Assume that $B$ is a smooth variety.
-There is a natural map, called the local Kodaira–Spencer map:
-(3.7) $k_{\pi_{B},b}^{\text{loc}}:T_{b}B\to\prod_{p\in({\mathcal{C}}_{b})_{\mathrm{sing}}}T\mathrm{Def}_{{\mathcal{C}}_{b},p}=H^{0}({\mathcal{C}}_{b},{\mathcal{E}xt}^{1}(\Omega_{{\mathcal{C}}_{b}},{\mathcal{O}}_{{\mathcal{C}}_{b}})),$
-where $({\mathcal{C}}_{b})_{\mathrm{sing}}$ is the singular locus of ${\mathcal{C}}_{b}\vcentcolon=\pi_{B}^{-1}(b)$, and $T\mathrm{Def}_{{\mathcal{C}}_{b},p}$ is the tangent space of the deformation functor of the complete local $\mathbb{C}$-algebra $\hat{{\mathcal{O}}}_{{\mathcal{C}}_{b},p}$.
+There is a natural map (cf.
+[13, Ch.
+II, Lemma 1.20, Exercise 1.4.4])
+(3.8) $\kappa_{\pi_{B},b}^{\text{loc}}:T_{b}B\to\prod_{p\in({\mathcal{C}}_{b})_{\mathrm{sing}}}T\mathrm{Def}_{{\mathcal{C}}_{b},p}=H^{0}({\mathcal{C}}_{b},{\mathcal{E}xt}^{1}(\Omega_{{\mathcal{C}}_{b}},{\mathcal{O}}_{{\mathcal{C}}_{b}})),$
+called the local Kodaira–Spencer map, where $({\mathcal{C}}_{b})_{\mathrm{sing}}$ is the singular locus of ${\mathcal{C}}_{b}\vcentcolon=\pi_{B}^{-1}(b)$, and $T\mathrm{Def}_{{\mathcal{C}}_{b},p}$ is the tangent space of the deformation functor of the complete local $\mathbb{C}$-algebra $\widehat{{\mathcal{O}}}_{{\mathcal{C}}_{b},p}$.
 Definition 3.4.
-The family $\pi_{B}:{\mathcal{C}}_{B}\to B$ is called locally versal if for any point $b\in B$, the local Kodaira–Spencer map $k_{\pi_{B},b}^{\text{loc}}$ in (3.7) is surjective.
-If $\pi_{B}:{\mathcal{C}}_{B}\to B$ is a family of integral planar curves, then for any $e\in\mathbb{Z}$, denote by
-$p_{B,e}:\overline{J}_{B}^{e}\to B$
-the relative compactified Jacobian of degree $e$ associated with $\pi_{B}$, whose fiber over a point $b\in B$ is the moduli space of rank $1$, torsion-free sheaves of degree $e$ (or equivalently, Euler characteristic $e+1-p_{a}({\mathcal{C}}_{b})$) on ${\mathcal{C}}_{b}$.
-Since $\pi_{B}$ admits, étale locally on $B$, a section which avoids the singularities of fibers of $\pi_{B}$, the smoothness of $\overline{J}_{B}^{e}$ for any $e$ is equivalent to the smoothness of $\overline{J}_{B}^{0}$.
-Hence, the following result follows directly from [7, Corollary B.3].
-Theorem 3.5 ([7]).
-If $\pi_{B}:{\mathcal{C}}_{B}\to B$ is a locally versal family of integral planar curves, then $\overline{J}_{B}^{e}$ is smooth for any $e\in\mathbb{Z}$.
+The family $\pi_{B}:{\mathcal{C}}_{B}\to B$ is called locally versal if for any point $b\in B$, the local Kodaira–Spencer map $\kappa_{\pi_{B},b}^{\text{loc}}$ in (3.8) is surjective.
+If $\pi_{B}:{\mathcal{C}}_{B}\to B$ is a family of integral, locally planar curves, then for any $e\in\mathbb{Z}$, denote by $\varrho_{B,e}:\overline{J}_{B}^{e}\to B$ the relative compactified Jacobian of degree $e$ associated with $\pi_{B}$.
+The fiber $\varrho_{B,e}^{-1}(b)$ over a point $b\in B$ is the moduli space of rank $1$, torsion-free sheaves of degree $e$ (or equivalently, with Euler characteristic $e+1-p_{a}({\mathcal{C}}_{b})$) on ${\mathcal{C}}_{b}$.
+Since $\pi_{B}$ admits, étale locally on $B$, a section which avoids the singularities of fibers of $\pi_{B}$, the smoothness of $\overline{J}_{B}^{e}$ is equivalent to that of $\overline{J}_{B}^{0}$.
+Hence, the following result follows directly from [10, Corollary B.3].
+Proposition 3.5 ([10, Corollary B.3]).
+If $\pi_{B}:{\mathcal{C}}_{B}\to B$ is a locally versal family of integral, locally planar curves, then $\overline{J}_{B}^{e}$ is smooth for any $e\in\mathbb{Z}$.
 3.2. Local versality and smoothness in high codimension
@@
 Proposition 3.6.
-Let $\ell\geq 2$ be an integer. If $\beta$ is $(2\ell-2)$-very ample, then there exists an open subscheme $V_{\ell}\subset|\beta|$ such that
+If $\beta$ is $\max\{1,2k-2\}$-very ample, then there exists an open subscheme $V_{k}\subset|\beta|$ such that
   1. (1)
-    Every point of $V_{\ell}$ corresponds to a reduced curve;
+    every point of $V_{k}$ corresponds to a reduced curve;
   2. (2)
-    $\operatorname{codim}(|\beta|\setminus V_{\ell},|\beta|)\geq\ell+1$;
+    $\operatorname{codim}(|\beta|\setminus V_{k},|\beta|)\geq k+1$;
   3. (3)
-    The universal family $\pi|_{\pi^{-1}(V_{\ell})}:\pi^{-1}(V_{\ell})\to V_{\ell}$ of curves in $V_{\ell}$ is locally versal.
+    the universal family $\pi|_{\pi^{-1}(V_{k})}:\pi^{-1}(V_{k})\to V_{k}$ of curves in $V_{k}$ is locally versal.
 Proof.
-We use an argument in [16] to control the locus of curves that are not $\ell$-nodal.
-Here, a curve is called $\ell$-nodal if it is reduced and has no singularities other than $\ell$ nodes.
-Since $\beta$ is also $\ell$-very ample, by the proof of [16, Proposition 2.1], there is an open subscheme $V_{\ell}\subset|\beta|$ such that
-  • $\operatorname{codim}(|\beta|\setminus V_{\ell},|\beta|)>\ell$.
-  • Curves in $V_{\ell}$ are either $\ell$-nodal, or reduced of geometric genus $p_{a}(\beta)-\ell$.
-Note that for a curve $C\in V_{\ell}$ which is not $\ell$-nodal, its geometric genus is
-(3.8) $p_{a}(\beta)-\sum_{p\in C_{\mathrm{sing}}}\delta_{p}(C)>p_{a}(\beta)-\ell.$
-By (3.8) and Milnor’s formula (3.6), we have
-$\sum_{p\in C_{\mathrm{sing}}}\tau_{p}(C)\leq\sum_{p\in C_{\mathrm{sing}}}\mu_{p}(C)\leq 2\sum_{p\in C_{\mathrm{sing}}}\delta_{p}(C)<2\ell.$
-It remains to show Property (3). For any curve $D\in V_{\ell}$, the map $k_{\pi,D}^{\text{loc}}$ in (3.7) can be identified with
-$\bar{r}_{W}:H^{0}({\mathcal{O}}_{S}(\beta))/\langle s_{D}\rangle\to H^{0}({\mathcal{O}}_{S}(\beta)|_{W})$
-induced by the restriction to a 0-dimensional closed subscheme $W\subset S$ of length
-$\sum_{p\in D_{\mathrm{sing}}}\tau_{p}(D)<2\ell$
-by Lemma 3.3, where $s_{D}\in H^{0}({\mathcal{O}}_{S}(\beta))$ is a defining section of $D$. Hence, $\bar{r}_{W}$ is surjective by the ($2\ell-2$)-very ampleness of $\beta$. ∎
-Combining the above proposition with Theorem 3.5, we can prove Theorem 1.2.
+By setting $V_{0}\vcentcolon=V_{1}$, we may assume $k>0$.
+The assumption on $\beta$ implies that $\beta$ is both $k$-very ample and (2k22k-2)-very ample.
+We use an argument in [22] to control the locus of curves that are not $k$-nodal.
+Here, a curve is called $k$-nodal if it is reduced and has no singularities other than $k$ nodes.
+Since $\beta$ is $k$-very ample, by the proof of [22, Proposition 2.1], there is an open subscheme $V_{k}\subset|\beta|$ satisfying the following conditions:
+  • $\operatorname{codim}(|\beta|\setminus V_{k},|\beta|)>k$.
+  • Curves in $V_{k}$ are either $k$-nodal, or reduced of geometric genus $p_{a}(\beta)-k$.
+Note that for a curve $C\in V_{k}$ which is not $k$-nodal, its geometric genus $p_{g}(C)$ satisfies
+$p_{g}(C)=p_{a}(\beta)-\sum_{p\in C_{\mathrm{sing}}}\delta_{p}(C)>p_{a}(\beta)-k.$
+From this and Milnor’s formula (3.7), we obtain
+(3.9) $\sum_{p\in C_{\mathrm{sing}}}\tau_{p}(C)\leq\sum_{p\in C_{\mathrm{sing}}}\mu_{p}(C)\leq 2\sum_{p\in C_{\mathrm{sing}}}\delta_{p}(C)<2k.$
+It remains to show Property (3).
+For any curve $D\in V_{k}$, let $s_{D}\in H^{0}(S,{\mathcal{O}}_{S}(\beta))$ be a defining section of $D$.
+The map $\kappa_{\pi,D}^{\text{loc}}$ as in (3.8) can be identified with the map
+$\bar{r}_{Z_{D}}:H^{0}(S,{\mathcal{O}}_{S}(\beta))/\langle s_{D}\rangle\to H^{0}(Z_{D},{\mathcal{O}}_{S}(\beta)|_{Z_{D}})$
+induced by restriction to a 0-dimensional closed subscheme $Z_{D}\subset S$.
+By (3.9) (which still holds for a $k$-nodal curve) and Lemma 3.3, the length $h^{0}({\mathcal{O}}_{Z_{D}})$ of $Z_{D}$ satisfies
+$h^{0}({\mathcal{O}}_{Z_{D}})=h^{0}({\mathcal{E}xt}^{1}(\Omega_{D},{\mathcal{O}}_{D}))=\sum_{p\in D_{\mathrm{sing}}}\tau_{p}(D)<2k.$
+Hence, $\kappa_{\pi,D}^{\text{loc}}$ is surjective by the (2k22k-2)-very ampleness of $\beta$. ∎
+Combining the above proposition with Proposition 3.5, we can prove Theorem 1.2.
 Proof of Theorem 1.2.
-Let $V_{\ell}$ be as in Proposition 3.6. Take $U_{\ell}=V_{\ell}\cap|\beta|^{\mathrm{int}}$. By Property (2) of $V_{\ell}$ in Proposition 3.6, we have
-$\operatorname{codim}(|\beta|^{\mathrm{int}}\setminus U_{\ell},|\beta|^{\mathrm{int}})\geq\ell+1.$
-It remains to show that $h^{-1}(U_{\ell})$ is smooth.
-Note that $h^{-1}(U_{\ell})$ is the relative compactified Jacobian of degree $p_{a}(\beta)-1+\chi$ associated with the universal family of curves in $U_{\ell}$.
-This universal family is locally versal by Property (3) of $V_{\ell}$ in Proposition 3.6.
-Thus, the smoothness of $h^{-1}(U_{\ell})$ follows from Theorem 3.5.
+Let $V_{k}$ be as in Proposition 3.6. Take $U_{k}=V_{k}\cap|\beta|^{\mathrm{int}}$. By Property (2) of $V_{k}$ in Proposition 3.6, we have
+$\operatorname{codim}(|\beta|^{\mathrm{int}}\setminus U_{k},|\beta|^{\mathrm{int}})\geq k+1.$
+It remains to show that $h^{-1}(U_{k})$ is smooth.
+Note that $h^{-1}(U_{k})$ is the relative compactified Jacobian of degree $p_{a}(\beta)-1+\chi$ associated with the universal family of curves in $U_{k}$.
+This universal family is locally versal by Property (3) of $V_{k}$ in Proposition 3.6.
+Thus, the smoothness of $h^{-1}(U_{k})$ follows from Proposition 3.5.
 ∎
 4. Intersection Betti numbers
-In this section, we generalize the approach in [29] to prove Theorem 1.1. For any open subscheme $U\subset|\beta|$, let
+In this section, we generalize the approach in [36] to prove Theorem 1.1.
+To start with, we review some properties of relative Hilbert schemes of points over $|\beta|$, which will play an important role in the proof.
+4.1. Relative Hilbert schemes of points
+Let $\pi:{\mathcal{C}}\to|\beta|$ be the universal family of curves in $|\beta|$. For $k\in\mathbb{Z}_{\geq 0}$, denote by
+$\pi^{[k]}:{\mathcal{C}}^{[k]}\to|\beta|$
+the relative Hilbert scheme of $k$ points on the fibers of $\pi$.
+For $C\in|\beta|$, the fiber of $\pi^{[k]}$ over $C$ is the Hilbert scheme $C^{[k]}$ parametrizing 0-dimensional, length $k$ closed subschemes of $C$.
+Note that $\pi^{[0]}:|\beta|\to|\beta|$ is the identity and $\pi=\pi^{[1]}:{\mathcal{C}}\to|\beta|$.
+The following property of $\pi^{[k]}$ is a corollary of [24, Theorem 1.1] (see also [42, Proposition 2.7]).
+Lemma 4.1.
+The morphism $\pi^{[k]}:{\mathcal{C}}^{[k]}\to|\beta|$ has fibers of the same dimension $k$.
+There is a natural morphism $\sigma_{k}:{\mathcal{C}}^{[k+1]}\to S^{[k+1]}$ defined by
+$Z\subset C]\mapsto[Z\subset S],$
+where $C\in|\beta|$, and $Z\subset C$ is a 0-dimensional closed subscheme of length $k+1$.
+The fiber of $\sigma_{k}$ over $Z\subset S]\in S^{[k+1]}$ is the projective space $\mathbb{P}(H^{0}(S,{\mathcal{I}}_{Z}\otimes{\mathcal{O}}_{S}(\beta))^{\vee})$, where ${\mathcal{I}}_{Z}$ is the ideal sheaf of $Z\subset S$.
+A well-known connection between $k$-very ampleness and the structure of ${\mathcal{C}}^{[k+1]}$ as a scheme over $S^{[k+1]}$ via $\sigma_{k}$ is as follows (e.g., [42, Proposition 2.6]).
+Lemma 4.2.
+If $\beta$ is $k$-very ample, then ${\mathcal{C}}^{[k+1]}$ is a projective bundle over $S^{[k+1]}$.
+Recall that for the Hilbert–Chow morphism $h:M_{\beta,\chi}\to|\beta|$, the fibers of $h$ over the locus $|\beta|^{\mathrm{int}}$ of integral curves are the compactified Jacobians of degree $p_{a}(\beta)-1+\chi$ and have dimension $p_{a}(\beta)$.
+For any open subscheme $U\subset|\beta|$, let $h_{U}:h^{-1}(U)\to U$ be the restriction of $h$, and write ${\mathcal{C}}_{U}^{[k]}$ for $(\pi^{[k]})^{-1}(U)$.
+Later, we will use the following result, due to Maulik–Yun [28, Theorem 2.13] and Migliorini–Shende [31, Corollary 2].
+Theorem 4.3 ([28, 31] [^2] The results of [28, 31] apply since the smoothness of $h^{-1}(U)$ implies, via a section of the universal curve (étale locally on $U$, avoiding the singularities of fibers), that the relative compactified Jacobians $\overline{J}_{U}^{e}$ for all $e\in\mathbb{Z}$ are smooth, and by [41, Proposition 14], so are the relative Hilbert schemes ${\mathcal{C}}^{[k]}_{U}$ for all $k\in\mathbb{Z}_{\geq 0}$.).
+Suppose that $|\beta|$ contains a connected smooth curve.
+Let $U\subset|\beta|^{\mathrm{int}}$ be an open subscheme such that $h^{-1}(U)$ is smooth, and let $P_{\bullet}H^{*}(h^{-1}(U))$ be the perverse filtration induced by $h_{U}:h^{-1}(U)\to U$.
+For any $k,\ell\in\mathbb{Z}_{\geq 0}$, we have an isomorphism of vector spaces
+(4.10) $H^{\ell}\left({\mathcal{C}}^{[k]}_{U}\right)\cong\bigoplus_{i+j\leq k;\,i,j\geq 0}\mathrm{Gr}_{i}^{P}H^{\ell-2j}\left(h^{-1}(U)\right).$
+Remark 4.4.
+When $k=0$, the isomorphism (4.10) reads
+$H^{\ell}(U)\cong\mathrm{Gr}_{0}^{P}H^{\ell}\left(h^{-1}(U)\right),$
+which can be obtained as a consequence of Ngô’s support theorem [33, Théorème 7.2.1] (see [28, Theorem 2.4, §2.5] or [31, Corollary 9]).
+Ngô’s support theorem implies that $h_{U}:h^{-1}(U)\to U$ has full support.
+More precisely, there is an isomorphism
+$Rh_{U*}{\mathbb{Q}}_{h^{-1}(U)}[\dim U]\simeq\bigoplus_{i=0}^{2p_{a}(\beta)}\operatorname{IC}(\wedge^{i}R^{1}\pi^{s}_{*}{\mathbb{Q}}_{{\mathcal{C}}_{U}^{\mathrm{sm}}})[-i],$
+where $\pi^{s}:{\mathcal{C}}_{U}^{\mathrm{sm}}\to U^{\mathrm{sm}}$ is the universal family of smooth curves in $U$.
+4.2. Strategy of the proof
+For any open subscheme $U\subset|\beta|$, let
 $N(U)\vcentcolon=2\operatorname{codim}(|\beta|\setminus U,|\beta|)-2.$
 We simply write $N(\beta)$ for $N(|\beta|^{\mathrm{int}})$.
-First, we focus on a smooth open subscheme $h^{-1}(U)$ of $M_{\beta,\chi}$ and prove the following result concerning the dimensions of graded pieces of the perverse filtration induced by the restriction of $h:M_{\beta,\chi}\to|\beta|$ (see (2.4)).
+As a preliminary step, we focus on a smooth open subscheme $h^{-1}(U)$ of $M_{\beta,\chi}$ and prove the following result concerning graded dimensions of the perverse filtration induced by the restriction $h_{U}:h^{-1}(U)\to U$ of $h:M_{\beta,\chi}\to|\beta|$ (see (2.6)).
 Theorem 4.5.
 Suppose that $|\beta|$ contains a connected smooth curve.
-Let $U\subset|\beta|^{\mathrm{int}}$ be an open subscheme such that $h^{-1}(U)$ is smooth, and let $h_{U}:h^{-1}(U)\to U$ be the restriction of $h$.
-For given $i,j\in\mathbb{Z}_{\geq 0}$, if $i+j\leq N(U)$, the divisor $\beta$ is (i1)(i-1)-very ample, and $2\dim|\beta|\geq 3i+j$, then
-(4.9) $n_{h_{U}}^{i,j}=n_{\infty}^{i,j},$
-where $n_{\infty}^{i,j}$ is the coefficient of $q^{i}t^{j}$ in the expansion of
+Let $U\subset|\beta|^{\mathrm{int}}$ be an open subscheme such that $h^{-1}(U)$ is smooth.
+For any $i,j\in\mathbb{Z}_{\geq 0}$ satisfying $i+j\leq N(U)$ and $3i+j\leq 2\dim|\beta|$, if $\beta$ is (i1)(i-1)-very ample, then
+$n_{h_{U}}^{i,j}=[H(q,t)]^{i,j},$
+where $H(q,t)]^{i,j}$ is the coefficient of $q^{i}t^{j}$ in the expansion of
 (4.11) $H(q,t)\vcentcolon=(1-qt)\prod_{m\geq 1}\frac{(1+q^{m}t^{m-1})^{b_{1}(S)}(1+q^{m}t^{m+1})^{b_{1}(S)}}{(1-q^{m+1}t^{m-1})(1-q^{m}t^{m})^{b_{2}(S)}(1-q^{m-1}t^{m+1})}.$
 Remark 4.6.
-Recall Göttsche’s formula [8] for $\sum_{n\geq 0}\sum_{i\geq 0}b_{i}(S^{[n]})z^{i}w^{n}$:
+Recall Göttsche’s formula [11] for $\sum_{k\geq 0}\sum_{i\geq 0}b_{i}(S^{[k]})z^{i}w^{k}$:
 (4.12) $G(z,w)\vcentcolon=\prod_{m\geq 1}\frac{(1+z^{2m-1}w^{m})^{b_{1}(S)}(1+z^{2m+1}w^{m})^{b_{1}(S)}}{(1-z^{2m-2}w^{m})(1-z^{2m}w^{m})^{b_{2}(S)}(1-z^{2m+2}w^{m})}.$
@@
 $z=t,\quad w=\frac{q}{t},$
-the product formula $H(q,t)$ in (4.10) satisfies the following equality
+the product formula $H(q,t)$ in (4.11) satisfies the following equality:
 $\frac{H(q,t)}{1-qt}=G(z,w)\cdot\frac{1-w}{1-z^{2}}.$
-By (2.4), the $k$-th Betti number of $h^{-1}(U)$ is
+By (2.6), the $k$-th Betti number of $h^{-1}(U)$ satisfies
 (4.13) $b_{k}(h^{-1}(U))=\sum_{i=0}^{k}n_{h_{U}}^{i,k-i}.$
@@
 Corollary 4.7.
-Let $k\in\mathbb{Z}_{\geq 0}$. In the setting of Theorem 4.1, if $k\leq\min\{N(U),2\dim|\beta|/3\}$ and $\beta$ is (k1k-1)-very ample, then
+In the setting of Theorem 4.5, if $k\leq\min\{N(U),2\dim|\beta|/3\}$ and $\beta$ is (k1k-1)-very ample, then
 $b_{k}(h^{-1}(U))=b_{k}^{\infty},$
 where $b^{\infty}_{k}$ is defined by (1.1).
-4.1. Strategy of the proof
-The proof of Theorem 4.1 consists of the following steps.
-Step 1. Relate $h^{-1}(U)$ to the relative Hilbert scheme ${\mathcal{C}}^{[n]}_{U}$. By (2.5), for any $\ell,m\in\mathbb{Z}_{\geq 0}$,
-(4.13) $b_{m}({\mathcal{C}}^{[\ell]}_{U})=\sum_{i+j\leq\ell;\,i,j\geq 0}n_{h_{U}}^{i,m-i-2j}.$
-Step 2. Calculate the Betti numbers of ${\mathcal{C}}^{[\ell]}_{U}$ using Proposition 2.3.
+We prove Theorem 4.5 by relating $h^{-1}(U)$ to the relative Hilbert schemes ${\mathcal{C}}^{[k]}_{U}$. By (4.10), for any $k,\ell\in\mathbb{Z}_{\geq 0}$,
+(4.14) $b_{\ell}({\mathcal{C}}^{[k]}_{U})=\sum_{i+j\leq k;\,i,j\geq 0}n_{h_{U}}^{i,\ell-i-2j}.$
+To calculate the Betti numbers of ${\mathcal{C}}^{[k]}_{U}$, we use Lemmas 4.1 and 4.2.
 Proposition 4.8.
-In the setting of Theorem 4.1, if $m\leq\min\{N(U),2\dim|\beta|-2\ell\}$ and $\beta$ is ($\ell-1$)-very ample, then
-(4.14) $b_{m}({\mathcal{C}}^{[\ell]}_{U})=\sum_{n=0}^{\lfloor m/2\rfloor}b_{m-2n}(S^{[\ell]}).$
+In the setting of Theorem 4.5, if $\ell\leq\min\{N(U),2\dim|\beta|-2k\}$ and $\beta$ is (k1k-1)-very ample, then
+(4.15) $b_{\ell}({\mathcal{C}}^{[k]}_{U})=\sum_{r=0}^{\lfloor\ell/2\rfloor}b_{\ell-2r}(S^{[k]}).$
 Proof.
-Since $\pi^{[\ell]}:{\mathcal{C}}^{[\ell]}\to|\beta|$ has fibers of dimension $\ell$ by Proposition 2.1, we have
-$\operatorname{codim}({\mathcal{C}}^{[\ell]}\setminus{\mathcal{C}}_{U}^{[\ell]},{\mathcal{C}}^{[\ell]})=\operatorname{codim}(|\beta|\setminus U,|\beta|).$
-By Proposition 2.3, the $(\ell-1)$-very ampleness of $\beta$ implies that $\mathcal{C}$$^{[\ell]}$ is a projective bundle over $S^{[\ell]}$.
-In particular, $\mathcal{C}$$^{[\ell]}$ is irreducible and smooth.
-Then it follows from [29, Lemma 2.4] (or the proof of Lemma 5.6) that for $m\leq N(U)$,
-$b_{m}({\mathcal{C}}^{[\ell]}_{U})=b_{m}({\mathcal{C}}^{[\ell]}).$
-Combining with
-$\dim{\mathcal{C}}^{[\ell]}-\dim S^{[\ell]}=\dim|\beta|-\ell\geq m/2$
-by our assumption, we conclude
-$b_{m}({\mathcal{C}}^{[\ell]}_{U})=b_{m}({\mathcal{C}}^{[\ell]})=b_{m}(S^{[\ell]}\times\mathbb{P}^{\dim|\beta|-\ell})=\sum_{n=0}^{\lfloor m/2\rfloor}b_{m-2n}(S^{[\ell]}),$
-which proves (4.14). ∎
-Since all Betti numbers of $S^{[\ell]}$ are determined by Göttsche’s formula (4.11), we can prove Theorem 4.1 by induction on $i$, combining (4.13) and (4.14).
+Since all fibers of $\pi^{[k]}:{\mathcal{C}}^{[k]}\to|\beta|$ have dimension $k$ by Lemma 4.1, we have
+$\operatorname{codim}({\mathcal{C}}^{[k]}\setminus{\mathcal{C}}_{U}^{[k]},{\mathcal{C}}^{[k]})=\operatorname{codim}(|\beta|\setminus U,|\beta|).$
+By Lemma 4.2, the (k1)(k-1)-very ampleness of $\beta$ implies that ${\mathcal{C}}^{[k]}$ is a projective bundle over $S^{[k]}$.
+In particular, ${\mathcal{C}}^{[k]}$ is irreducible and smooth.
+Then it follows from [36, Lemma 2.4] (or the proof of Lemma 5.3) that for $\ell\leq N(U)$,
+$b_{\ell}({\mathcal{C}}^{[k]}_{U})=b_{\ell}({\mathcal{C}}^{[k]}).$
+Combining this with
+$\dim{\mathcal{C}}^{[k]}-\dim S^{[k]}=\dim|\beta|-k\geq\ell/2,$
+which follows from our assumption, we conclude that
+$b_{\ell}({\mathcal{C}}^{[k]}_{U})=b_{\ell}({\mathcal{C}}^{[k]})=b_{\ell}(S^{[k]}\times\mathbb{P}^{\dim|\beta|-k})=\sum_{r=0}^{\lfloor\ell/2\rfloor}b_{\ell-2r}(S^{[k]}),$
+thereby completing the proof of (4.15). ∎
+Since all Betti numbers of $S^{[k]}$ are determined by Göttsche’s formula (4.12), now we can prove Theorem 4.5 by combining (4.14) and (4.15).
 Proof of Theorem 4.5.
-We proceed by induction on $i$ to prove (4.9). For $i=0$ and $j\leq\min\{N(U),2\dim|\beta|\}$,
-$n_{h_{U}}^{0,j}$ $=\dim P_{0}H^{j}(h^{-1}(U))$
+We use induction on $i$. For $i=0$ and $j\leq\min\{N(U),2\dim|\beta|\}$,
+$n_{h_{U}}^{0,j}$ $=\dim_{\mathbb{Q}}\mathrm{Gr}_{0}^{P}H^{j}(h^{-1}(U))$
 […]
@@
 […]
-which proves the induction base of (4.9).
-Suppose the result is true for $(i^{\prime},j^{\prime})$ with $i^{\prime}<i$ and $i^{\prime}+j^{\prime}\leq i+j$.
-For $m\leq i+j$, it follows by (4.13), our induction hypothesis, and (4.14) that
-$n_{h_{U}}^{i,m-i}$ $=\sum_{i^{\prime}+j^{\prime}=i;\,i^{\prime},j^{\prime}\geq 0}n_{h_{U}}^{i^{\prime},m-i^{\prime}-2j^{\prime}}-\sum_{i^{\prime}<i;\,i^{\prime}+j^{\prime}=i;\,i^{\prime},j^{\prime}\geq 0}n_{h_{U}}^{i^{\prime},m-i^{\prime}-2j^{\prime}}$
-$=\left(b_{m}({\mathcal{C}}^{[i]}_{U})-b_{m}({\mathcal{C}}^{[i-1]}_{U})\right)-\sum_{i^{\prime}<i;\,i^{\prime}+j^{\prime}=i;\,i^{\prime},j^{\prime}\geq 0}n_{\infty}^{i^{\prime},m-i^{\prime}-2j^{\prime}}$
-$=\sum_{n=0}^{\lfloor m/2\rfloor}\left(b_{m-2n}(S^{[i]})-b_{m-2n}(S^{[i-1]})\right)-\sum_{i^{\prime}<i;\,i^{\prime}+j^{\prime}=i;\,i^{\prime},j^{\prime}\geq 0}n_{\infty}^{i^{\prime},m-i^{\prime}-2j^{\prime}}$
-$=n_{\infty}^{i,m-i},$
-where the last identity is essentially the same as [29, (31)] by Göttsche’s formula [8] (see also Remark 4.2). ∎
-Finally, we prove our first main theorem stated in the introduction.
+which proves the induction base.
+Then suppose that $n_{h_{U}}^{i^{\prime},j^{\prime}}=[H(q,t)]^{i^{\prime},j^{\prime}}$ holds for any $i^{\prime},j^{\prime}\in\mathbb{Z}_{\geq 0}$ satisfying $i^{\prime}<i$ and $i^{\prime}+j^{\prime}\leq i+j$.
+For any $\ell\leq i+j$, it follows from (4.14), our induction hypothesis, and (4.15) that
+$n_{h_{U}}^{i,\ell-i}$ $=\sum_{i^{\prime}+j^{\prime}=i;\,i^{\prime},j^{\prime}\geq 0}n_{h_{U}}^{i^{\prime},\ell-i^{\prime}-2j^{\prime}}-\sum_{i^{\prime}<i;\,i^{\prime}+j^{\prime}=i;\,i^{\prime},j^{\prime}\geq 0}n_{h_{U}}^{i^{\prime},\ell-i^{\prime}-2j^{\prime}}$
+$=\left(b_{\ell}({\mathcal{C}}^{[i]}_{U})-b_{\ell}({\mathcal{C}}^{[i-1]}_{U})\right)-\sum_{i^{\prime}<i;\,i^{\prime}+j^{\prime}=i;\,i^{\prime},j^{\prime}\geq 0}[H(q,t)]^{i^{\prime},\ell-i^{\prime}-2j^{\prime}}$
+$=\sum_{r=0}^{\lfloor\ell/2\rfloor}\left(b_{\ell-2r}(S^{[i]})-b_{\ell-2r}(S^{[i-1]})\right)-\sum_{i^{\prime}<i;\,i^{\prime}+j^{\prime}=i;\,i^{\prime},j^{\prime}\geq 0}[H(q,t)]^{i^{\prime},\ell-i^{\prime}-2j^{\prime}}$
+$=[H(q,t)]^{i,\ell-i},$
+where the last identity is essentially equivalent to [36, (31)] and follows from Remark 4.6. This completes the induction step and hence the proof. ∎
+Finally, we prove our first main theorem stated in Section 1.
 Proof of Theorem 1.1.
-By the upper semicontinuity of fiber dimensions, we have
+By the upper semicontinuity of fiber dimension, we have
 (4.16) $N(\beta)\geq 2c-2.$
-Since $\beta$ is $\max\{2,2k-2\}$-very ample, let $U_{k}\subset|\beta|^{\mathrm{int}}$ be as in Theorem 1.2 (set $U_{k}\vcentcolon=U_{2}$ if $k\leq 1$).
+Since $\beta$ is $\max\{1,2k-2\}$-very ample, let $U_{k}\subset|\beta|^{\mathrm{int}}$ be as in Theorem 1.2.
 Using (4.16), the assumption $k\leq c-1$, and Theorem 1.2, we obtain the following properties:
+  1. (1)
     $N(U_{k})\geq\min\{N(\beta),2k\}=2k$;
+  2. (2)
     $h^{-1}(U_{k})$ is smooth.
-By Corollary 4.3, we have
+Note that $|\beta|$ contains a connected smooth curve by Bertini’s theorem. Applying Corollary 4.7 with $U=U_{k}$, we have
 (4.17) $b_{k}(h^{-1}(U_{k}))=b_{k}^{\infty}.$
@@
 $\operatorname{codim}(M_{\beta,\chi}\setminus h^{-1}(U_{k}),M_{\beta,\chi})\geq\min\{c,k+1\}>k.$
-Then by [23, Theorem 6.7.4] and the smoothness of $h^{-1}(U_{k})$, we have isomorphisms
-$\operatorname{IH}^{k}(M_{\beta,\chi})\cong\operatorname{IH}^{k}(h^{-1}(U_{k}))\cong H^{k}(h^{-1}(U_{k})),$
-and therefore
-$\dim_{\mathbb{Q}}\operatorname{IH}^{k}(M_{\beta,\chi})=b_{k}(h^{-1}(U_{k}))=b_{k}^{\infty}$
-by (4.16), which completes the proof. ∎
+Then by [30, Theorem 6.7.4] and the smoothness of $h^{-1}(U_{k})$, we have isomorphisms
+$I\!H^{k}(M_{\beta,\chi})\cong I\!H^{k}(h^{-1}(U_{k}))\cong H^{k}(h^{-1}(U_{k})),$
+and therefore, by (4.17),
+$\dim_{\mathbb{Q}}I\!H^{k}(M_{\beta,\chi})=b_{k}(h^{-1}(U_{k}))=b_{k}^{\infty},$
+which completes the proof. ∎
 4.3. Discussion on the general case
-In general, the moduli space $M_{\beta,\chi}$ can be reducible even if $\beta$ is sufficiently positive; see [34, §5].
-Moreover, even if we assume that $M_{\beta,\chi}$ is irreducible, it is usually hard to compute the codimension $c$ in Theorem 1.1.
-It is not necessarily true that $c$ is the same as $\operatorname{codim}(|\beta|\setminus|\beta|^{\mathrm{int}},|\beta|)$.
-In some nice cases where $c=\operatorname{codim}(|\beta|\setminus|\beta|^{\mathrm{int}},|\beta|)$, the asymptotic behavior of $\operatorname{IH}^{k}(M_{\beta,\chi})$ can be described more precisely, provided that $\operatorname{codim}(|\beta|\setminus|\beta|^{\mathrm{int}},|\beta|)$, or equivalently $N(\beta)$, can be calculated.
+In general, the moduli space $M_{\beta,\chi}$ can be reducible even if $\beta$ is sufficiently positive (see [42, §5]).
+Moreover, even if we assume that $M_{\beta,\chi}$ is irreducible, it might not be true that the codimension $c$ in Theorem 1.1 is the same as $\operatorname{codim}(|\beta|\setminus|\beta|^{\mathrm{int}},|\beta|)$.
+In the cases where $c=\operatorname{codim}(|\beta|\setminus|\beta|^{\mathrm{int}},|\beta|)$, the asymptotic behavior of $I\!H^{k}(M_{\beta,\chi})$ can be described more precisely, provided that $\operatorname{codim}(|\beta|\setminus|\beta|^{\mathrm{int}},|\beta|)$, or equivalently $N(\beta)$ (defined in §4.2), can be calculated.
 The following lemma allows us to estimate $N(\beta)$ by analyzing the decompositions of $\beta$.
@@
 Proof.
-Let $D_{1},\cdots,D_{\rho}$ be an orthogonal basis for $\operatorname{Num}(S)_{\mathbb{Q}}$, where $D_{1}$ is the class of an ample divisor.
-By Hodge index theorem, $D_{\ell}^{2}<0$ for $1<\ell\leq\rho$.
+Let $D_{1},\cdots,D_{\rho}$ be an orthogonal basis for $\operatorname{Num}(S)_{\mathbb{Q}}\vcentcolon=\operatorname{Num}(S)\otimes_{\mathbb{Z}}{\mathbb{Q}}$, where $D_{1}$ is the class of an ample divisor.
+By the Hodge index theorem, $D_{\ell}^{2}<0$ for $1<\ell\leq\rho$.
 Given any $C_{1}$, $C_{2}$ as in the statement, write
 $\beta\equiv\sum_{j=1}^{\rho}a_{j}D_{j},\quad C_{i}\equiv\sum_{j=1}^{\rho}a_{i,j}D_{j}$
-for some $a_{j},a_{i,j}\in{\mathbb{Q}}$ ($i=1,2$), where $\equiv$ denotes the numerical equivalence.
-For $1<\ell\leq\rho$, let $n_{\ell}>0$ be the integer such that $A_{\ell}\vcentcolon=n_{\ell}D_{1}+D_{\ell}$ is ample.
+for some $a_{j},a_{i,j}\in{\mathbb{Q}}$, where $\equiv$ denotes the numerical equivalence.
+For $1<\ell\leq\rho$, let $n_{\ell}>0$ be an integer such that $A_{\ell}\vcentcolon=n_{\ell}D_{1}+D_{\ell}$ is ample.
 Then $D_{1}\cdot C_{i}>0$ and $A_{\ell}\cdot C_{i}>0$, which implies
@@
 Combined with $D_{\ell}^{2}<0$ and $C_{1}+C_{2}=\beta$, this shows that the possible values of $a_{i,j}$ ($i=1,2$; $j=1,\cdots,\rho$) are bounded since $\beta$ is fixed, and thus the result follows. ∎
-By the calculations in Section 5, we ask the following question.
-Question 4.6.
+Motivated by the del Pezzo case ([36, Proposition 1.5]), along with the cases studied in Section 5, we ask the following question.
+Question 4.10.
 For any ample divisor $\beta$ on an arbitrary smooth projective surface $S$, is it true that
 $\lim_{d\to+\infty}N(d\beta)=+\infty?$
-5. Stabilization in the Enriques and bielliptic cases
-In this section, let $S$ be an Enriques or bielliptic surface.
+5. Stabilization for minimal surfaces of Kodaira dimension zero
+Throughout this section, except in §5.4, we consider a minimal surface $S$ of Kodaira dimension 0.
 In particular, its canonical divisor $K_{S}$ is numerically trivial.
-When $\beta$ is a base-point-free, ample divisor on $S$, $|\beta|$ contains a connected smooth curve by Bertini’s theorem.
-Denote by $|\beta|^{\mathrm{sm}}\subset|\beta|$ the open subscheme parametrizing smooth curves.
-To apply Theorem 1.1, we first prove the irreducibility of the moduli spaces of one-dimensional sheaves on these surfaces.
+In this situation, the following proposition gives an affirmative answer to Question 4.10, which we then use to prove Theorem 1.4.
 Proposition 5.1.
-Suppose that $S$ is an Enriques or bielliptic surface, and $\beta$ is a base-point-free, ample divisor satisfying $\operatorname{gcd}(\beta\cdot H,\chi)=1$. Then $M_{\beta,\chi}$ is irreducible.
-Proof.
-Since the open subscheme $h^{-1}(|\beta|^{\mathrm{int}})$ is irreducible and contains the smooth open subscheme $h^{-1}(|\beta|^{\mathrm{sm}})$ by [34, Proposition 2.2], there is a unique irreducible component $Y\subset M_{\beta,\chi}$ containing $h^{-1}(|\beta|^{\mathrm{int}})$.
-We prove the irreducibility of $M_{\beta,\chi}$ by contradiction.
-If $Y\not=M_{\beta,\chi}$, then take another irreducible component $Y^{\prime}\neq Y$.
-It follows from $Y^{\prime}\cap h^{-1}(|\beta|^{\mathrm{sm}})=\emptyset$ that $Y^{\prime}\subset h^{-1}(|\beta|\setminus|\beta|^{\mathrm{sm}})$.
-By [36, Corollary 1.3] and the fact that $K_{S}$ is numerically trivial, the fibers of $h:M_{\beta,\chi}\to|\beta|$ have the same dimension $p_{a}(\beta)$; therefore
-(5.17) $\dim Y^{\prime}\leq\dim h^{-1}(|\beta|\setminus|\beta|^{\mathrm{sm}})<\dim|\beta|+p_{a}(\beta)=\beta^{2}+\chi({\mathcal{O}}_{S}),$
-where the last equality follows from the Kodaira vanishing theorem and the Riemann–Roch formula.
-However, using the morphism $\operatorname{det}:M\to\operatorname{Pic}(X)$ in [11, Theorem 4.5.4]  [^2] Although [11, Theorem 4.5.4] is stated for moduli spaces of positive rank sheaves, the proof of $\dim_{[F]}M({\mathcal{Q}})\geq\dim_{[F]}M-\dim_{[\operatorname{det}F]}\operatorname{Pic}(X)$ does not rely on the positive rank assumption.
-and the Riemann–Roch formula, for any point ${\mathcal{F}}\in Y^{\prime}$, we have
-$\dim Y^{\prime}$ $\geq\operatorname{ext}^{1}({\mathcal{F}},{\mathcal{F}})-\operatorname{ext}^{2}({\mathcal{F}},{\mathcal{F}})-h^{1}({\mathcal{O}}_{S})$
-$=\beta^{2}+\chi({\mathcal{O}}_{S}),$
-which contradicts (5.17). ∎
-Then we assume the following property of $N(\beta)$, whose proof will be given in §5.1.2 and §5.2.2 by analyzing the Enriques and bielliptic cases separately.
-Proposition 5.2.
-If $\beta$ is an ample divisor on an Enriques or bielliptic surface, then
+If $\beta$ is an ample divisor on a minimal surface $S$ of Kodaira dimension 0, then
 $\lim_{d\to+\infty}N(d\beta)=+\infty.$
-Combining the above results with Theorem 1.1, we can prove the stabilization of intersection Betti numbers of $M_{\beta}=M_{\beta,1}$.
-Theorem 5.3.
-Suppose that $S$ is an Enriques or bielliptic surface.
-Fix $k\in\mathbb{Z}_{\geq 0}$.
-Given any ample divisor $\beta_{0}$ on $S$, there exists $d(\beta_{0},k)\in\mathbb{Z}_{>0}$ (depending on $\beta_{0}$ and $k$) such that
-(5.18) $\dim_{\mathbb{Q}}\operatorname{IH}^{k}(M_{d\beta_{0}})=b_{k}^{\infty}\quad\text{for all integer }d\geq d(\beta_{0},k),$
-where $b_{k}^{\infty}$ is defined by (1.1).
-Proof.
+The proof of Proposition 5.1 will be given in §5.1.2, §5.2.2, and §5.3, where we analyze the Enriques, bielliptic, K3 and Abelian cases separately.
+Together with Proposition 2.2 and Proposition 5.1, Theorem 1.1 implies the stabilization of the intersection Betti numbers of $M_{\beta,\chi}$.
+Proof of Theorem 1.4.
 Choose $d(\beta_{0},k)\in\mathbb{Z}_{>0}$ such that for all $d\geq d(\beta_{0},k)$, the following properties hold:
@@
   2. (2)
-    $d\beta_{0}$ is $\max\{2,2k-2\}$-very ample;
+    $d\beta_{0}$ is $\max\{1,2k-2\}$-very ample;
   3. (3)
-    $2\dim|d\beta_{0}|\geq 3k$.
-Since the fibers of $h:M_{d\beta_{0}}\to|d\beta_{0}|$ have the same dimension by [36, Corollary 1.3], we have
-(5.19) $\operatorname{codim}(M_{d\beta_{0}}\setminus h^{-1}(|d\beta_{0}|^{\mathrm{int}}),M_{d\beta_{0}})$ $=\operatorname{codim}(|d\beta_{0}|\setminus|d\beta_{0}|^{\mathrm{int}},|d\beta_{0}|)$
+    $2\dim|d\beta_{0}|=2\chi({\mathcal{O}}_{S}(d\beta_{0}))-2\geq 3k$.
+By Proposition 2.2 (i), $M_{d\beta_{0},\chi}$ is irreducible, and by Proposition 2.2 (ii), we have
+$\operatorname{codim}(M_{d\beta_{0},\chi}\setminus h^{-1}(|d\beta_{0}|^{\mathrm{int}}),M_{d\beta_{0},\chi})$ $=\operatorname{codim}(|d\beta_{0}|\setminus|d\beta_{0}|^{\mathrm{int}},|d\beta_{0}|)$
 $=1+\frac{1}{2}N(d\beta_{0})\geq k+1.$
-Hence, the result follows from Proposition 5.1 and Theorem 1.1 by taking $\beta=d\beta_{0}$. ∎
-Remark 5.4.
-When $\beta\cdot H$ and $\chi$ are not necessarily coprime, there can be strictly semistable sheaves in $M_{\beta,\chi}$ and thus $h:M_{\beta,\chi}\to|\beta|$ may not have constant fiber dimension by [36].
-In this case, one could possibly prove the irreducibility of $M_{\beta,\chi}$ by showing that the corresponding moduli stack is irreducible (cf.
-[19, §2.6]), but the first equality of (5.19) may fail.
-Therefore, to ensure the equality in (5.18) holds for $M_{d\beta_{0},\chi}$, we need an additional assumption that $d\beta_{0}\cdot H$ is coprime with $\chi$.
+Hence, the result follows from Theorem 1.1 with $\beta=d\beta_{0}$. ∎
 5.1. The case of Enriques surfaces
@@
 Its second Betti number is $b_{2}(S)=10$.
-The following property of linear systems on $S$ is well-known (e.g., [13, Theorem]).
+The following property of linear systems on $S$ is well-known (e.g., [19, Theorem]) and was used in [37, Proposition 1.14] to compute the codimension of the locus of reducible curves in a linear system.
 Lemma 5.2.
-Let $D$ be a nonzero, nef, effective divisor on an Enriques surface. We have the following properties:
+Let $D$ be a nonzero effective nef divisor on an Enriques surface $S$.
   1. (i)
@@
   2. (ii)
-    If $D^{2}=0$, then $D$ is numerically equivalent to $kE$ for some $k\in\mathbb{Z}_{>0}$ and a primitive, effective divisor $E$ with $E^{2}=0$. Moreover,
-    $\dim|D|=\begin{cases}\lfloor\frac{k}{2}\rfloor&(\text{if }D=kE)\\ \lfloor\frac{k-1}{2}\rfloor&(\text{if }D=kE+K_{S}).\end{cases}$
+    If $D^{2}=0$, then $D$ is linearly equivalent to either $mE$ or $mE+K_{S}$ for some $m\in\mathbb{Z}_{>0}$ and some primitive effective divisor $E$ with $E^{2}=0$. Moreover,
+    $\dim|D|=\begin{cases}\lfloor\frac{m}{2}\rfloor&(\text{if }D\sim mE)\\ \lfloor\frac{m-1}{2}\rfloor&(\text{if }D\sim mE+K_{S}).\end{cases}$
 5.1.2. Proof of Proposition 5.1 in the Enriques case
-Any non-integral curves in $|d\beta|$ are in the image of $|C_{1}|\times|C_{2}|$ for finitely many decompositions $d\beta=C_{1}+C_{2}$ (up to linear or numerical equivalence) by Lemma 4.5, where $C_{1}$ and $C_{2}$ are nonzero effective divisors on $S$.
+By Lemma 4.9, there are finitely many pairs of nonzero effective divisors $(C_{1},C_{2})$ (up to linear or numerical equivalence) with $d\beta=C_{1}+C_{2}$.
+Any non-integral curve in $|d\beta|$ lies in the image of $|C_{1}|\times|C_{2}|$ for one of these pairs.
 First we assume that $C_{1}$ and $C_{2}$ have nef irreducible components. To prove the lemma, we estimate all possible values of
@@
 (5.18) $\dim|d\beta|-\dim(|C_{1}|\times|C_{2}|)=C_{1}\cdot C_{2}\geq d\sqrt{2\beta^{2}}-2.$
-Case 1.2: $C_{1}^{2}=0$ and $C_{2}^{2}>0$. By Lemma 5.5 (ii), $C_{1}=sE_{1}$ or $sE_{1}+K_{S}$ for some $s\in\mathbb{Z}_{>0}$ and a primitive effective divisor $E_{1}$ with $E_{1}^{2}=0$. Moreover,
-(5.21) $\dim|d\beta|-\dim(|C_{1}|\times|C_{2}|)$ $\geq C_{1}\cdot C_{2}-\lfloor\frac{s}{2}\rfloor$
-$\geq s\left(d\beta\cdot E_{1}-\frac{1}{2}\right)\geq d-\frac{1}{2}.$
+Case 1.2: $C_{1}^{2}=0$ and $C_{2}^{2}>0$.
+By Lemma 5.2 (ii), $C_{1}$ is linearly equivalent to $m_{1}E_{1}$ or $m_{1}E_{1}+K_{S}$ for some $m_{1}\in\mathbb{Z}_{>0}$ and some primitive effective divisor $E_{1}$ with $E_{1}^{2}=0$.
+Moreover,
+(5.19) $\dim|d\beta|-\dim(|C_{1}|\times|C_{2}|)$ $\geq C_{1}\cdot C_{2}-\lfloor\frac{m_{1}}{2}\rfloor$
+$\geq m_{1}\left(d\beta\cdot E_{1}-\frac{1}{2}\right)\geq d-\frac{1}{2}.$
 Case 1.3: $C_{1}^{2}=C_{2}^{2}=0$.
-Let $s$ and $E_{1}$ be as in Case 1.2.
-By Lemma 5.5 (ii), $C_{2}=tE_{2}$ or $tE_{2}+K_{S}$ for some $t\in\mathbb{Z}_{>0}$ and a primitive effective divisor $E_{2}$ with $E_{2}^{2}=0$.
-Then we have $C_{1}\cdot C_{2}=stE_{1}\cdot E_{2}=d^{2}\beta^{2}/2>0$ and
-(5.22) $\dim|d\beta|-\dim(|C_{1}|\times|C_{2}|)$ $\geq C_{1}\cdot C_{2}-\lfloor\frac{s}{2}\rfloor-\lfloor\frac{t}{2}\rfloor\geq\frac{d^{2}\beta^{2}-(s+t)}{2}$
+Let $m_{1}$ and $E_{1}$ be as in Case 1.2.
+By Lemma 5.2 (ii), $C_{2}$ is linearly equivalent to $m_{2}E_{2}$ or $m_{2}E_{2}+K_{S}$ for some $m_{2}\in\mathbb{Z}_{>0}$ and some primitive effective divisor $E_{2}$ with $E_{2}^{2}=0$.
+Then we have $C_{1}\cdot C_{2}=m_{1}m_{2}E_{1}\cdot E_{2}=d^{2}\beta^{2}/2>0$ and
+(5.20) $\dim|d\beta|-\dim(|C_{1}|\times|C_{2}|)$ $\geq C_{1}\cdot C_{2}-\lfloor\frac{m_{1}}{2}\rfloor-\lfloor\frac{m_{2}}{2}\rfloor\geq\frac{d^{2}\beta^{2}-(m_{1}+m_{2})}{2}$
 $\geq\frac{d^{2}\beta^{2}-(1+d^{2}\beta^{2}/2)}{2}=\frac{d^{2}\beta^{2}-2}{4}.$
@@
 (5.21) $\dim|d\beta|-\dim|C_{1}|=\frac{d\beta\cdot C_{2}+C_{1}\cdot C_{2}}{2}\geq\frac{d}{2}.$
-Case 2.2: $C_{1}^{2}=0$. Let $s$ and $E_{1}$ be as in Case 1.2. By Lemma 5.5 (ii),
-(5.24) $\dim|d\beta|-\dim|C_{1}|\geq\frac{d^{2}\beta^{2}}{2}-\lfloor\frac{s}{2}\rfloor\geq\frac{s(d\beta\cdot E_{1}-1)+d\beta\cdot C_{2}}{2}\geq d-\frac{1}{2}.$
+Case 2.2: $C_{1}^{2}=0$. Let $m_{1}$ and $E_{1}$ be as in Case 1.2. By Lemma 5.2 (ii),
+(5.22) $\dim|d\beta|-\dim|C_{1}|\geq\frac{d^{2}\beta^{2}}{2}-\lfloor\frac{m_{1}}{2}\rfloor\geq\frac{m_{1}(d\beta\cdot E_{1}-1)+d\beta\cdot C_{2}}{2}\geq d-\frac{1}{2}.$
 Combining (5.18)–(5.22) concludes the proof. ∎
 5.1.3. Refined invariants
-Assume that $S$ is a generic Enriques surface and that the class of $\beta$ in $\operatorname{Num}(S)$ is not divisible by $2$ so that the moduli space $M_{\beta}=M_{\beta,1}$ is a smooth projective Calabi–Yau variety by [30].
-Denote by $n_{\beta}^{i,j}$ the number defined by (2.4) with $\phi$ taken to be $h:M_{\beta}\to|\beta|$.
-In order to apply Theorem 4.1 to compute $n_{\beta}^{i,j}$, we need the following lemma which is parallel to [20, Corollary 5.3]. For the reader’s convenience, we provide a proof below.
+Assume that $S$ is a generic Enriques surface.
+For an ample divisor $\beta$ whose class in $\operatorname{Num}(S)$ is not divisible by $2$, the moduli space $M_{\beta}=M_{\beta,1}$ is a smooth projective Calabi–Yau variety by [37].
+The linear system $|\beta|$ contains a connected smooth curve (e.g., [37, Corollary 1.13]).
+Denote by $n_{\beta}^{i,j}$ the number defined by (2.6) with $\varphi$ taken to be $h:M_{\beta}\to|\beta|$.
+In order to apply Theorem 4.5 to compute $n_{\beta}^{i,j}$, we need the following lemma which is parallel to [26, Corollary 5.3]. For the reader’s convenience, we provide a proof below.
 Lemma 5.3.
-Let $\beta$ be a base-point-free, ample divisor on a generic Enriques surface $S$ such that $2\nmid\beta$ in $\operatorname{Num}(S)$. When $i+j\leq N(\beta)$, we have
+Let $\beta$ be an ample divisor on a generic Enriques surface $S$ with $2\nmid\beta$ in $\operatorname{Num}(S)$. For any $i,j\in\mathbb{Z}_{\geq 0}$ satisfying $i+j\leq N(\beta)$, we have
 $n_{\beta}^{i,j}=n^{i,j}_{h^{\circ}},$
-where $h^{\circ}=h_{|\beta|^{\mathrm{int}}}:h^{-1}(|\beta|^{\mathrm{int}})\to|\beta|^{\mathrm{int}}$.
+where $h^{\circ}=h_{|\beta|^{\mathrm{int}}}:h^{-1}(|\beta|^{\mathrm{int}})\to|\beta|^{\mathrm{int}}$ is the restriction of $h:M_{\beta}\to|\beta|$.
 Proof.
 Let $W=M_{\beta}\setminus h^{-1}(|\beta|^{\mathrm{int}})$.
-Denote by $\iota:W\hookrightarrow M_{\beta}$ the closed embedding, and by $\jmath:h^{-1}(|\beta|^{\mathrm{int}})\hookrightarrow M_{\beta}$, $\jmath^{\circ}:|\beta|^{\mathrm{int}}\to|\beta|$ the open embeddings.
-Since the fibers of $h$ have the same dimension, we have
+Denote by $\iota:W\hookrightarrow M_{\beta}$ the closed embedding, and by $\jmath:h^{-1}(|\beta|^{\mathrm{int}})\hookrightarrow M_{\beta}$, $\jmath^{\circ}:|\beta|^{\mathrm{int}}\hookrightarrow|\beta|$ the open embeddings.
+By Proposition 2.2 (ii),
 $\operatorname{codim}(W,M_{\beta})=\operatorname{codim}(|\beta|\setminus|\beta|^{\mathrm{int}},|\beta|)=1+\frac{1}{2}N(\beta).$
-Consider the exact triangle
-$\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}}\longrightarrow{\mathbb{Q}}_{M_{\beta}}\longrightarrow\jmath_{*}\jmath^{*}{\mathbb{Q}}_{M_{\beta}}\stackrel{{\scriptstyle+1}}{{\longrightarrow}}$
-which yields another exact triangle
-$h_{*}\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}}\longrightarrow h_{*}{\mathbb{Q}}_{M_{\beta}}\longrightarrow\jmath^{\circ}_{*}h^{\circ}_{*}{\mathbb{Q}}_{h^{-1}(|\beta|^{\mathrm{int}})}\stackrel{{\scriptstyle+1}}{{\longrightarrow}}.$
-By the property of Verdier duals ([23, Proposition 5.3.9]) and the smoothness of $M_{\beta}$,
-$\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}}\cong(D_{M_{\beta}}\iota_{*}{\mathbb{Q}}_{W})[-2\dim M_{\beta}],$
+Consider the exact triangle (all functors are derived)
+$\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}}\longrightarrow{\mathbb{Q}}_{M_{\beta}}\longrightarrow\jmath_{*}\jmath^{*}{\mathbb{Q}}_{M_{\beta}}\stackrel{{\scriptstyle+1}}{{\longrightarrow}}.$
+By the property of Verdier duals [30, Proposition 5.3.9] and the smoothness of $M_{\beta}$,
+$\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}}\simeq(D_{M_{\beta}}\iota_{*}{\mathbb{Q}}_{W})[-2\dim M_{\beta}],$
 and therefore for $k\leq N(\beta)+1$,
-$H^{k}(\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}})=H^{2\dim M_{\beta}-k}(W)^{\vee}=0.$
-By taking the long exact sequence of cohomologies, the vanishing of $H^{k}(\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}})$ shows that for $\ell\leq N(\beta)$, the restriction map induces an isomorphism
+${\mathbb{H}}^{k}(M_{\beta},\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}})\cong H^{2\dim M_{\beta}-k}(W)^{\vee}=0.$
+By the long exact sequence in hypercohomology, the vanishing of ${\mathbb{H}}^{k}(M_{\beta},\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}})$ shows that for $\ell\leq N(\beta)$, the restriction map induces an isomorphism
 (5.23) $H^{\ell}(M_{\beta})\cong H^{\ell}(h^{-1}(|\beta|^{\mathrm{int}})).$
-Since the restriction to $h^{-1}(|\beta|^{\mathrm{int}})$ clearly preserves a relatively ample class (relative to $h$), it follows from [6, Lemma 3.3] that $n_{\beta}^{i,j}\leq n_{h^{\circ}}^{i,j}$ for $i+j\leq N(\beta)$.
+Since the restriction to $h^{-1}(|\beta|^{\mathrm{int}})$ clearly preserves a relatively ample class (relative to $h$), it follows from [7, Lemma 3.3] that $n_{\beta}^{i,j}\leq n_{h^{\circ}}^{i,j}$ for $i+j\leq N(\beta)$.
 Then the result follows by (4.13) and (5.23).
 ∎
-Now we can prove the stabilization of $n_{\beta}^{i,j}$ for the moduli space of one-dimensional sheaves on a generic Enriques surface.
-Theorem 5.7.
-Fix $i,j\in\mathbb{Z}_{\geq 0}$.
-Given any ample divisor $\beta_{0}$ on a generic Enriques surface $S$ such that $2\nmid\beta_{0}$ in $\operatorname{Num}(S)$, there exists $d(\beta_{0},i,j)\in\mathbb{Z}_{>0}$ (depending on $\beta_{0}$, $i$ and $j$) such that
-(5.26) $n_{d\beta_{0}}^{i,j}=n_{\infty}^{i,j}\quad\text{for all integer $d\geq d(\beta_{0},i,j)$ with }2\nmid d,$
-where $n_{\infty}^{i,j}$ is determined by
-$\sum_{i,j\geq 0}n_{\infty}^{i,j}q^{i}t^{j}=(1-qt)\prod_{m\geq 1}\frac{1}{(1-q^{m+1}t^{m-1})(1-q^{m}t^{m})^{10}(1-q^{m-1}t^{m+1})}.$
-Proof.
-Choose the integer $d(\beta_{0},i,j)\in\mathbb{Z}_{>0}$ such that for all $d\geq d(\beta_{0},i,j)$, the following properties hold:
+Now we can prove the stabilization of $n_{\beta}^{i,j}$ for moduli spaces of one-dimensional sheaves on a generic Enriques surface.
+Proof of Theorem 1.6.
+Choose $d(\beta_{0},i,j)$ such that for all $d\geq d(\beta_{0},i,j)$, the following properties hold:
   1. (1)
@@
   2. (2)
-    $d\beta_{0}$ is $\max\{0,i-1\}$-very ample;
+    $d\beta_{0}$ is (i1)(i-1)-very ample;
   3. (3)
     $d^{2}\beta_{0}^{2}\geq 3i+j$.
-Condition (3) is equivalent to $2\dim|d\beta_{0}|\geq 3i+j$ by Lemma 5.5 (i). By Lemma 5.6 and Theorem 4.1, taking $U=|d\beta_{0}|^{\mathrm{int}}$, we have
-$n_{d\beta_{0}}^{i,j}=n_{h_{U}}^{i,j}=n_{\infty}^{i,j},$
-which proves (5.26). ∎
+Property (3) is equivalent to $2\dim|d\beta_{0}|\geq 3i+j$ by Lemma 5.2 (i). Applying Lemma 5.3 and Theorem 4.5 with $\beta=d\beta_{0}$ and $U=|d\beta_{0}|^{\mathrm{int}}$, we have
+$n_{d\beta_{0}}^{i,j}=n_{h_{U}}^{i,j}=[H(q,t)]^{i,j}=n_{\infty}^{i,j},$
+which completes the proof. ∎
 Again by (4.13), the next corollary follows immediately.
 Corollary 5.4.
-Fix $k\in\mathbb{Z}_{\geq 0}$.
-Given any ample divisor $\beta_{0}$ on a generic Enriques surface $S$ such that $2\nmid\beta_{0}$ in $\operatorname{Num}(S)$, there exists $d(\beta_{0},k)\in\mathbb{Z}_{>0}$ (depending on $\beta_{0}$ and $k$) such that
+Given any ample divisor $\beta_{0}$ on a generic Enriques surface $S$ with $2\nmid\beta_{0}$ in $\operatorname{Num}(S)$ and any $k\in\mathbb{Z}_{\geq 0}$, there exists $d(\beta_{0},k)\in\mathbb{Z}_{>0}$ such that
 $b_{k}(M_{d\beta_{0}})=b_{k}^{\infty}\quad\text{for all $d\geq d(\beta_{0},k)$ with }2\nmid d,$
-where $b_{k}^{\infty}$ is determined by (1.1).
-Remark 5.9.
-By (5.20)–(5.22) and [35, Proposition 2.5], we can take
-(5.27) $d(\beta_{0},i,j)=\max\left\{2,\,i+1,\,\left\lceil\frac{i+j+2}{2}\right\rceil,\,\left\lceil\frac{i+j+6}{2\sqrt{2\beta_{0}^{2}}}\right\rceil,\,\left\lceil\sqrt{\frac{2i+2j+6}{\beta_{0}^{2}}}\,\right\rceil\right\}$
-in Theorem 1.4 (Theorem 5.7).
+where $b_{k}^{\infty}$ is defined by (1.1).
+Remark $5.$
+Note that a generic Enriques surface does not contain any (2)(-2)-curves. By (5.18)–(5.20) and [44, Proposition 2.5], we can take
+(5.24) $d(\beta_{0},i,j)=\max\left\{i+1,\,\left\lceil\frac{i+j+2}{2}\right\rceil,\,\left\lceil\frac{i+j+6}{2\sqrt{2\beta_{0}^{2}}}\right\rceil,\,\left\lceil\sqrt{\frac{2i+2j+6}{\beta_{0}^{2}}}\,\right\rceil\right\}$
+in Theorem 1.6.
 5.2. The case of bielliptic surfaces
 5.2.1. Basic properties
-Recall that a bielliptic surface $S$ is a minimal surface of Kodaira dimension 0 with $h^{1}({\mathcal{O}}_{S})=1$ and $h^{2}({\mathcal{O}}_{S})=0$. In particular, the exponential sequence reads
+Recall that a bielliptic surface $S$ is a minimal surface of Kodaira dimension 0 with $h^{1}({\mathcal{O}}_{S})=1$ and $h^{2}({\mathcal{O}}_{S})=0$. In this case, the exponential sequence reads
 $0\to\operatorname{Pic}^{0}(S)\to\operatorname{Pic}(S)\to H^{2}(S,\mathbb{Z})\to 0,$
-where $\operatorname{Pic}^{0}(S)=H^{1}({\mathcal{O}}_{S})/H^{1}(S,\mathbb{Z})$ is the group of algebraically trivial divisors modulo linear equivalence and has the structure of an elliptic curve.
-There exist two elliptic curves $A$, $B$, and an abelian group $G$ acting on $A$ and $B$, such that ([31, Proposition 1.1])
+where $\operatorname{Pic}^{0}(S)=H^{1}(S,{\mathcal{O}}_{S})/H^{1}(S,\mathbb{Z})$ is the group of algebraically trivial divisors modulo linear equivalence and has the structure of an elliptic curve.
+By [1, pp.
+112–114] (see also [39, Proposition 1.1]), there exist two elliptic curves $A$, $B$, together with a finite Abelian group $G$ acting on $A$ and $B$, such that
   1. (1)
-    $A/G$ is elliptic and $B/G\cong\mathbb{P}^{1}$;
+    $A/G$ is an elliptic curve and $B/G\cong\mathbb{P}^{1}$;
   2. (2)
-    $S\cong(A\times B)/G$, where $G$ acts on $A\times B$ componentwise;
-see [1].
-We follow the convention in [31] by abuse of notation to denote by $A$ (resp.
-$B$) the class (in $\operatorname{Num}(S)$, $H^{2}(S,\mathbb{Z})$, or $H^{2}(S,{\mathbb{Q}})$) of a smooth fiber of the natural projection $\Psi:S\to B/G$ (resp.
-$\Phi:S\to A/G$).
-These classes satisfy
+    $S\cong(A\times B)/G$, where $G$ acts on $A\times B$ componentwise.
+We follow the convention in [39] to denote also by $A$ (resp. $B$) the class in $\operatorname{Num}(S)$ of a smooth fiber of the natural projection $\Psi:S\to B/G$ (resp. $\Phi:S\to A/G$). Then
 $A^{2}=B^{2}=0,\quad A\cdot B=\gamma,$
-where $\gamma$ is the order of $G$. Let $\lambda A$, $\mu B$ be a basis for $\operatorname{Num}(S)$ ($\lambda,\mu\in{\mathbb{Q}}$); see [31, Theorem 1.4] for the values of $\lambda$ and $\mu$.
-We need the following lemma to compute the dimension of linear systems.
-Lemma 5.10 ([31, Lemma 1.3]).
-Let $D$ be a divisor of numerical class $sA+tB$ with $s,t\in{\mathbb{Q}}$ on a bielliptic surface. The following properties hold:
+where $\gamma$ is the order of $G$.
+There exist $\lambda,\mu\in{\mathbb{Q}}_{>0}$ such that the classes $\lambda A$ and $\mu B$ form a basis of $\operatorname{Num}(S)$ (see [39, Theorem 1.4] for the values of $\lambda$ and $\mu$).
+We need the following lemma to compute the dimension of linear systems on $S$.
+Lemma 5.6 ([39, Lemma 1.3]).
+Let $D$ be a divisor of numerical class $uA+vB$ with $u,v\in{\mathbb{Q}}$ on a bielliptic surface $S$. The following properties hold:
   1. (i)
-    $\chi(\mathcal{O}_{S}(D))=st\gamma$.
+    $\chi(\mathcal{O}_{S}(D))=uv\gamma$.
   2. (ii)
-    $D$ is ample if and only if s>0,t>0s>0,\,t>0.
+    $D$ is ample if and only if u>0,v>0u>0,\,v>0.
   3. (iii)
@@
   4. (iv)
-    If $H^{0}(\mathcal{O}_{S}(D))\neq 0$, then $s\geq 0,\,t\geq 0$.
+    If $h^{0}(\mathcal{O}_{S}(D))\neq 0$, then $u\geq 0,\,v\geq 0$.
 5.2.2. Proof of Proposition 5.1 in the bielliptic case
-The numerical class of $\beta$ in $\operatorname{Num}(S)$ is of the form $a\lambda A+b\mu B$ with $a,b\in\mathbb{Z}_{>0}$. By Lemma 5.10 (i) and (iii),
+By Lemma 5.6 (i)–(iii), the numerical class of $\beta$ in $\operatorname{Num}(S)$ is of the form $a\lambda A+b\mu B$ with $a,b\in\mathbb{Z}_{>0}$, and
 $\dim|d\beta|=\chi({\mathcal{O}}_{S}(d\beta))-1=d^{2}ab\lambda\mu\gamma-1.$
-Any non-integral curve in $|d\beta|$ is in the image of a morphism from a scheme $P$ to $|d\beta|$, where $P$ parametrizes the triples $(C,D,N)$ satisfying $N\in\operatorname{Pic}^{0}(S)$, $C\in|C_{1}+N|$, and $D\in|C_{2}-N|$ for two nonzero effective divisors $C_{1}$, $C_{2}$ such that $C_{1}+C_{2}=d\beta$.
+Any non-integral curve in $|d\beta|$ is in the image of a morphism from a scheme $P$ to $|d\beta|$, where $P$ parametrizes the triples $(C,D,L)$ satisfying $L\in\operatorname{Pic}^{0}(S)$, $C\in|C_{1}+L|$, and $D\in|C_{2}-L|$ for two nonzero effective divisors $C_{1}$, $C_{2}$ such that $C_{1}+C_{2}=d\beta$.
 By Lemma 4.9 (or Lemma 5.6), the images of finitely many such morphisms cover $|d\beta|\setminus|d\beta|^{\mathrm{int}}$.
@@
 $\dim|C_{1}|+\dim|C_{2}|=\chi({\mathcal{O}}_{S}(C_{1}))+\chi({\mathcal{O}}_{S}(C_{2}))-2.$
-Therefore, by the proof of [29, (13)] (or Case 1.1 in §5.1.2), we have
-(5.28) $\dim|d\beta|-\dim|C_{1}|-\dim|C_{2}|-1=C_{1}\cdot C_{2}\geq d\sqrt{\beta^{2}}-1.$
+Therefore, by the proof of (5.18), we have
+(5.25) $\dim|d\beta|-\dim|C_{1}|-\dim|C_{2}|-1=C_{1}\cdot C_{2}\geq d\sqrt{2\beta^{2}}-2.$
 Case 2: At least one of $a_{1},a_{2},b_{1},b_{2}$ is zero. We may assume $a_{1}=0$. Then $b_{1}>0$ since otherwise $C_{1}$ would be numerically trivial. By Lemma 5.6 (i)–(iii),
 (5.26) $\dim|C_{1}|\leq\dim|A+C_{1}|=b_{1}\mu\gamma-1.$
-If $b_{2}>0$, then it follows from $b_{1}>0$ and (5.29) that
-(5.30) $\dim|d\beta|-\dim|C_{1}|-\dim|C_{2}|-1$ $=C_{1}\cdot C_{2}-\dim|C_{1}|$
-$\geq dab_{1}\lambda\mu\gamma-(b_{1}\mu\gamma-1)$
-$\geq(da\lambda-1)\mu\gamma+1\quad(\text{if }da\lambda>1).$
-If $b_{2}=0$, then
+If $b_{2}>0$, then it follows from Lemma 5.6 (i)–(iii), (5.26), and $b_{1}>0$ that
+(5.27) $\dim|d\beta|-\dim|C_{1}|-\dim|C_{2}|-1$ $=C_{1}\cdot C_{2}-\dim|C_{1}|-1$
+$\geq dab_{1}\lambda\mu\gamma-(b_{1}\mu\gamma-1)-1$
+$\geq(da\lambda-1)\mu\gamma\quad(\text{if }da\lambda\geq 1).$
+If $b_{2}=0$, then as in (5.26), we have
 $\dim|C_{2}|\leq\dim|C_{2}+B|=da\lambda\gamma-1,$
@@
 Combining (5.25), (5.27), and (5.28) completes the proof. ∎
+5.3. The case of K3 or Abelian surfaces
+Now let $S$ be a K3 or an Abelian surface.
+The proof of Proposition 5.1 in this case is very similar to the proofs for Enriques and bielliptic surfaces (given in §5.1.2 and §5.2.2, respectively), so we keep it brief.
+Since the Kawamata–Viehweg vanishing theorem shows that Lemma 5.2 (i) is also valid for K3 and Abelian surfaces, we only need the following lemma (see [19, Theorem] for the K3 case, and [17, Proposition 2.3, Corollary 2.6] for the Abelian case).
+Lemma 5.7.
+Let $D$ be a nonzero effective nef divisor on a K3 or an Abelian surface $S$ with $D^{2}=0$.
+Then $D$ is numerically equivalent to $mE$ for some $m\in\mathbb{Z}_{>0}$ and some primitive effective divisor $E$ with $E^{2}=0$.
+Moreover,
+[…]
+5.4. Some related stabilization questions
+We collect some stabilization questions for moduli spaces of one-dimensional sheaves.
+Now $S$ is not necessarily a minimal surface of Kodaira dimension 0.
+For simplicity, we may assume $h^{1}({\mathcal{O}}_{S})=0$.
+We begin by providing a refined version of Conjecture 1.3, stated heuristically as follows.
+Conjecture 5.8.
+For fixed $i,j\in\mathbb{Z}_{\geq 0}$ and $\chi\in\mathbb{Z}$, the perverse Hodge number $n_{h}^{i,j}$ of $h:M_{\beta,\chi}\rightarrow|\beta|$ stabilizes as $\beta$ becomes sufficiently positive.
+Furthermore, $n_{h}^{i,j}$ stabilizes to $H(q,t)]^{i,j}$ (defined as in Theorem 4.5) if $M_{\beta,\chi}$ is asymptotically irreducible.
+In [36], the approach to Conjecture 1.3 for del Pezzo surfaces is to prove Conjecture 5.8.
+When $S$ is a K3 or an Abelian surface and $M_{\beta,\chi}$ is smooth, Shen–Yin [40] proved that the perverse Hodge numbers of $M_{\beta,\chi}$ are equal to its Hodge numbers using hyper-Kähler geometry and thus the stabilization of the perverse Hodge numbers follows from the closed formula of Hodge numbers of $S^{[n]}$ [12, Theorem 2.3.14].
+Our Theorem 1.6 also shows the stabilization of perverse Hodge numbers when $S$ is an Enriques surface and $M_{\beta,\chi}$ is smooth.
+Question 5.9.
+For fixed $i,j\in\mathbb{Z}_{\geq 0}$ and $\chi\in\mathbb{Z}$, does the intersection Hodge number $\dim_{\mathbb{Q}}I\!H^{i,j}(M_{\beta,\chi})$ stabilize as $\beta$ becomes sufficiently positive?
+It seems that our method in this paper does not work for intersection Hodge numbers in the cases of Enriques and bielliptic surfaces.
+In [29], Mauri and Migliorini showed that Question 5.9 has a positive answer for K3 surfaces, based on the previously known case where $\chi=1$ and $S$ is a K3 surface.
+The stabilization problems for moduli spaces of one-dimensional sheaves on $S$ can also be studied from a motivic perspective.
+Let $K_{0}(\mathrm{Var})$ be the Grothendieck ring of varieties, which is the free Abelian group generated by (the isomorphism classes of) varieties $[X]$ modulo the scissor relations
+$[X]=[Y]+[Z],$
+where $Y$ and $Z$ are disjoint locally closed subvarieties of $X$ with $X=Y\cup Z$.
+The product in $K_{0}(\mathrm{Var})$ is defined by taking the fiber product.
+Let $\mathbb{L}\in K_{0}(\mathrm{Var})$ be the class of the affine line $\mathbb{A}^{1}$.
+Denote by $\widehat{K}_{0}(\mathrm{Var})$ the completion of $K_{0}(\mathrm{Var})[\mathbb{L}^{-1}]$ with respect to the dimension filtration (see [8, p.
+17] for more details).
+Question 5.10.
+For fixed $\chi\in\mathbb{Z}$, does the class $M_{\beta,\chi}]$ stabilize in $\widehat{K}_{0}(\mathrm{Var})$ as $\beta$ becomes sufficiently positive?
+Coskun and Woolf [8] proved that when $S$ is a rational surface and $K_{S}\cdot H<0$, the class of the moduli space of $H$-stable torsion-free sheaves with fixed Chern character on $S$ stabilizes in $\widehat{K}_{0}(\mathrm{Var})$ as the determinant of sheaves tends to $\infty$.
+A positive answer to Question 5.10 would establish an analogous result in the rank 0 case.
+A natural further question is whether there is a geometric interpretation for the stabilization of Betti numbers.
+One potential approach is to consider the tautological classes in $H^{*}(M_{\beta,\chi})$, i.e., the Künneth components of Chern classes of the universal sheaf (assuming its existence) on $M_{\beta,\chi}\times S$.
+Let $RH^{*}(M_{\beta,\chi})\subset H^{*}(M_{\beta,\chi})$ be the subring generated by the tautological classes.
+Although $RH^{*}(M_{\beta,\chi})\neq H^{*}(M_{\beta,\chi})$ in general (cf.
+[37, Theorem 5.10]), we expect the stabilization of Betti numbers of $M_{\beta,\chi}$ to be explained by the stabilization of $RH^{*}(M_{\beta,\chi})$.
+When $S$ is a del Pezzo surface and $M_{\beta,\chi}$ is smooth, one has $RH^{*}(M_{\beta,\chi})=H^{*}(M_{\beta,\chi})$ and the stable Betti numbers are determined by the combinatorics of tautological classes (see [36, Corollary 1.8]).
+If our expectation is true, the stabilization phenomenon in the del Pezzo case fits into a more general picture.
+As for the intersection cohomology, the $k$-th stable intersection Betti number calculated in this paper also coincides with $b_{k}^{\infty}$.
+In what follows, we assume that $M_{\beta,\chi}$ is pure dimensional.
+Note that there is a homomorphism of graded vector spaces (e.g., [30, Exercise 6.7.1, Remark 6.7.2])
+$\alpha:H^{*}(M_{\beta,\chi})\rightarrow I\!H^{*}(M_{\beta,\chi}).$
+Thus we can define $RI\!H^{\ast}(M_{\beta,\chi})$ as the image of $RH^{\ast}(M_{\beta,\chi})$ under the map $\alpha$.
+Given our results, we also expect the stabilization of intersection Betti numbers to be explained by the stabilization of ${RI\!H}^{\ast}(M_{\beta,\chi})$, and therefore we ask the following question.
+Question 5.11.
+Is it true that $\dim_{\mathbb{Q}}{RI\!H}^{k}(M_{\beta,\chi})=b_{k}^{\infty}$ if $\beta$ is sufficiently positive?
+Another direction is to explore the tautological relations in $RH^{\ast}(M_{\beta,\chi})$.
+When $S=\mathbb{P}^{2}$, the tautological relations are studied in [20], leading to the formulation of a strengthened version of the “P=C” conjecture proposed in [21].
+For other surfaces, however, the problem remains unexplored.
+More generally, one may consider similar questions for moduli spaces of one-dimensional sheaves on a projective surface with mild singularities.
+In this setting, we continue to use the notation $M_{\beta,\chi}$.
+Question 5.12.
+Fix $\chi\in\mathbb{Z}$ and let $\beta$ be a divisor on a simply connected projective surface with ADE singularities.
+For fixed $k\in\mathbb{Z}_{\geq 0}$, does the intersection Betti number $\dim_{\mathbb{Q}}I\!H^{k}(M_{\beta,\chi})$ stabilize as $\beta$ becomes sufficiently positive?
+Similarly, for fixed $i,j\in\mathbb{Z}_{\geq 0}$, does the intersection Hodge number $\dim_{\mathbb{Q}}I\!H^{i,j}(M_{\beta,\chi})$ stabilize?
+Toward answering Question 5.12, we expect our strategy in this paper to work in some cases after suitable modifications, but there are some difficulties.
+For example, the relative Hilbert schemes of points over the linear system $|\beta|$ may not be equidimensional as the curves in $|\beta|$ may not be locally planar.
+Moreover, the Betti numbers (or intersection Betti numbers) of Hilbert schemes of points on a singular surface are still unknown.
+Acknowledgments
+The authors would like to thank Mirko Mauri, Weite Pi, and Junliang Shen for helpful discussions.
+FZ is grateful to her advisor Jun Li for his constant support, to János Kollár and Claire Voisin for inspiring discussions on the deformation theory of curve singularities, to Georg Oberdieck for suggesting considering the case of bielliptic surfaces, and to Songsong Huang for useful conversations.
+FS is partially supported by the Fundamental Research Funds for the Central Universities, NSFC (No.
+12025106) and Shaanxi NSF (No.
+2025JC-QYCX-002).
+FZ is supported by the NSFC grants (No.
+12121001 and No.
+12425105) and LMNS.